### v1

# spacing=nonfrench

Abstract
Large Language Models (LLMs) have advanced code generation and software automation, but are fundamentally constrained by limited inference-time context and lack of explicit code structure reasoning. We introduce Kodezi Chronos [51], a next-generation architecture for autonomous code understanding, debugging, and maintenance, designed to operate across ultra-long contexts comprising entire codebases, histories, and documentation—all without fixed window limits.
Kodezi Chronos leverages a multi-level embedding memory engine, combining vector and graph-based indexing with continuous code-aware retrieval. This enables efficient and accurate reasoning over millions of lines of code, supporting repository-scale comprehension, multi-file refactoring, and real-time self-healing actions.
Our evaluation introduces a novel Multi Random Retrieval benchmark, specifically tailored to the software engineering domain. Unlike classical retrieval benchmarks, this method requires the model to resolve arbitrarily distant and obfuscated associations across code artifacts, simulating realistic tasks such as variable tracing, dependency migration, and semantic bug localization. Chronos outperforms prior LLMs and code models—demonstrating a 23% improvement in real-world bug detection and reducing debugging cycles by up to 40% compared to traditional sequence-based approaches.
By natively interfacing with IDEs and CI/CD workflows, Chronos enables seamless, autonomous software maintenance, elevating code reliability and productivity while reducing manual effort. These results mark a critical advance toward self-sustaining, continuously optimized software ecosystems.
I INTRODUCTION
Recent advancements in large language models (LLMs) have transformed code generation, review, and reasoning tasks [1, 2, 10]. However, debugging—the most time-consuming and critical aspect of software development—remains largely unsolved. While tools like GitHub Copilot [11], Cursor, and Claude excel at code completion [12], they fundamentally misunderstand debugging as a multi-faceted, context-heavy process that spans entire repositories, historical commits, CI/CD logs, and runtime behaviors. Production debugging requires reasoning across files separated by thousands of lines [13], understanding temporal code evolution, and correlating seemingly unrelated symptoms to root causes buried deep in dependency chains [14].
Current code assistants fail at debugging for three critical reasons: (1) they are trained primarily on code completion tasks, not debugging workflows [15]; (2) they lack persistent memory of past bugs, fixes, and codebase-specific patterns [16]; and (3) their context windows, even when extended to 100K+ tokens, cannot capture the full debugging context needed for complex, multi-file issues. Recent studies show that even state-of-the-art models like GPT-4, Claude-3, and Gemini-1.5 achieve less than 10% success rates on real-world debugging benchmarks [17, 18], often proposing superficial fixes that fail validation or introduce new regressions [19].
Kodezi Chronos represents a paradigm shift: the first debugging language model developed by Kodezi [51], specifically designed, trained, and optimized for autonomous bug detection, root cause analysis, and validated fix generation. Unlike code completion models that generate syntactically correct but often semantically flawed suggestions, Chronos operates through a continuous debugging loop—proposing fixes, running tests, analyzing failures, and iteratively refining solutions until validation succeeds. Built on a novel architecture combining persistent debug memory, multi-source retrieval (code, logs, traces, PRs), and execution sandboxing, Chronos achieves what no existing model can: true autonomous debugging at repository scale.
Chronos is designed for seamless integration with modern development stacks: it operates as an embedded “AI CTO” within CI/CD pipelines, IDEs, and project management tools. Its proactive, event-driven workflow ensures that debugging, documentation generation, refactoring, and even preventative maintenance actions occur autonomously, guided by deep repository memory rather than manual prompts or brittle heuristics.
To rigorously evaluate Chronos’s unique capabilities, we move beyond traditional benchmarks and propose a multi-step, random retrieval evaluation reflecting the authentic complexities of code search, dependency resolution, and semantic bug localization at scale. Empirically, Chronos demonstrates state-of-the-art results across industry-standard metrics and realistic maintenance tasks, reducing debugging times and increasing project resilience.
This paper presents the architecture, memory system, retrieval mechanism, evaluation methodology, and experimental results that establish Kodezi Chronos as the new frontier for autonomous, ultra-contextual codebase management.
II RELATED WORK AND LIMITATIONS OF TRADITIONAL APPROACHES
The emergence of large-scale neural models for source code processing—such as CodeBERT [46], GraphCodeBERT [4], and CodeT5 [5]—has significantly advanced automatic code synthesis, translation, and review. Many of these models, as well as massive LLMs like GPT-3 and GPT-4 [1, 47], are pre-trained on billions of lines of code paired with natural language, learning rich semantic representations for many programming languages.
Despite impressive gains on benchmark tasks, these approaches are fundamentally bottlenecked by attention-based architectures and fixed-size input windows, typically constraining context to tens of thousands of tokens. Practical developer workflows, by contrast, require tools to reason across entire repositories: files, modules, historic commits, documentation, and evolving build systems, far outstripping standard context limits.
State-of-the-art retrieval-augmented generation (RAG) methods extend practical context by embedding and recalling external documents [8], but remain mostly limited to chunked passage retrieval and lack true compositional, dependency-aware memory. Window expansion techniques—such as models with 100k tokens or more [7]—incur prohibitive compute/memory costs and suffer from diluted attention, leading to information loss and degraded performance as codebase size increases.
Existing benchmarks for long-context reasoning, like the ”Needle in a Haystack” pattern, commonly involve embedding a salient but easily differentiable clue in a sea of noise or repetition. Such evaluations unintentionally rely on the model’s ability to match explicit tokens or unusual patterns, providing only a partial test of true retrieval and reasoning ability. In real-world maintenance, however, code elements are frequently unmarked, distributed, and semantically interdependent, requiring the AI not just to ”find,” but to infer context, resolve relationships, and reason about impact [45].
Other recent work explores graph neural networks (GNNs) for modeling explicit data flow or control flow graphs [9, 6]. While GNNs can capture structural properties of code, they are rarely coupled with ultra-long context LLMs, and lack the continuous learning, memory updating, and rapid recall required for live autonomous maintenance.
Kodezi Chronos is motivated by these challenges: By combining continuous graph-aware indexing, dynamic embedding updates, and reasoning-optimized memory retrieval, Chronos transcends traditional limitations and enables truly repository-scale, real-time software comprehension and intervention.
III ARCHITECTURE OF KODEZI CHRONOS
III-A The Output-Heavy Nature of Debugging
Despite the industry focus on ever-larger context windows (128K, 200K, 1M+ tokens), debugging presents a fundamentally different challenge: it is inherently output-heavy rather than input-heavy. This asymmetry has profound implications for model design and optimization.
III-A1 Input vs Output Token Distribution
What models typically see (input):
  • Error stack traces: 200-500 tokens
  • Relevant source code: 1K-4K tokens
  • Test failures and logs: 500-2K tokens
  • Prior fix attempts: 500-1K tokens
  • Total input: Often ¡10K tokens for most real-world debugging tasks
What models must produce (output):
  • Multi-file bug fixes: 500-1,500 tokens
  • Root cause explanations: 300-600 tokens
  • Updated unit tests: 400-800 tokens
  • Commit messages/PR summaries: 150-300 tokens
  • Documentation updates: 200-400 tokens
  • Total output: Typically 2,000-4,000 tokens per debugging session
III-A2 Why Output Quality Trumps Input Size
TABLE I: Input vs output characteristics in debugging tasks.
Aspect Input Context Output Generation
Nature Sparse, localized Dense, structured
Cost Impact Sublinear with retrieval Linear to exponential
Quality Limiter Retrieval precision Generation accuracy
Success Factor Context relevance Syntactic & semantic correctness
The critical insight: a model with intelligent 8K context that generates robust, test-passing fixes will outperform a 1M-context model that produces syntactically correct but semantically flawed patches.
III-A3 Chronos’s Output-Optimized Architecture
Chronos addresses this asymmetry through several architectural innovations:
  1. 1.
    Debug-Specific Generation Training: Unlike code completion models trained on next-token prediction, Chronos is trained on complete debugging sessions, learning to generate structured fixes, explanations, and tests as cohesive units.
  2. 2.
    Iterative Refinement Loop: Rather than single-shot generation, Chronos validates outputs through execution, using test results to refine patches—ensuring output quality over quantity.
  3. 3.
    Template-Aware Generation: Chronos learns repository-specific patterns for commits, tests, and documentation, reducing output token waste while maintaining consistency.
  4. 4.
    Confidence-Guided Output: The model generates explanations and fallback strategies only when confidence is below threshold, optimizing output token usage.
This output-centric design enables Chronos to achieve 65.3% debugging success despite competitors having 10-100x larger context windows, validating that for debugging, output quality and structure matter more than input capacity.
III-B Core Architecture Overview
Kodezi Chronos is designed as an autonomous memory-driven intelligence layer for code, operating at scales that span entire enterprise repositories, team histories, and auxiliary knowledge sources. Its architecture consists of three core modules: (i) a persistent Memory Engine for continuous graph-based context construction, (ii) an advanced Retriever that dynamically assembles deep semantic context from code and documentation, and (iii) a transformer-based Code Reasoning Model for synthesis, debugging, and orchestration of software changes.
Code, Docs, CI/CD LogsMemory Engine(Embedding + Graph)Multi-CodeAssociation RetrieverReasoning Model& OrchestrationTest ResultsPatches, Changelogs,Test Results
Figure 1: High-level overview of Chronos: Memory-driven embedding and retrieval powering autonomous reasoning and codebase management.
[…]
Figure 2: Token distribution in debugging tasks: Unlike typical LLM applications where input dominates, debugging requires substantial, high-quality output generation.
III-C Memory Engine: Repository-Scale Embedding and Indexing
The Memory Engine is responsible for ingesting, encoding, and maintaining a unified semantic representation of all project files, versioned code, documentation, configuration, historical diffs, test outcomes, and architectural artifacts. Each unit of code (e.g., function level, file, commit) is parsed to extract both sequence and structural cues—using ASTs, dependency graphs, and metadata links.
These elements are then projected into a high-dimensional vector space via context-aware, code-specific encoders trained on multiple programming languages and coding paradigms. The Memory Engine stores not just static embeddings but also maintains an evolving graph database, where nodes represent code elements and edges denote code relationships (e.g., function calls, module imports, bug-ticket links, commit ancestry).
login.pyauth_utils.pyREADME.mdtest_login.pybug_1234.mdcallscited inlinked bug intested by
Figure 3: Graph-structured memory indexing in Kodezi Chronos: code, documentation, and test elements as nodes, with functional relationships as edges.
This design enables Chronos to efficiently retrieve, traverse, and reason about segments of the codebase that share non-local relationships, even if separated by thousands of lines, multiple files, or extensive revision history.
III-D Achieving Unlimited Context Through Smart Retrieval
Traditional LLMs are fundamentally constrained by attention complexity and memory limitations. Even models claiming ”unlimited” context achieve this through sliding windows or hierarchical attention that loses critical debugging information. Chronos implements true unlimited context through:
  • Hierarchical Code Embeddings: Multi-level representations from token → statement → function → module → repository
  • Temporal Context Indexing: Every code element tagged with commit history, allowing time-travel debugging
  • Semantic Dependency Graphs: Explicit modeling of import chains, inheritance hierarchies, and data flows
  • Dynamic Context Assembly: At inference, retrieves precisely the code paths relevant to the current bug
This approach enables Chronos to maintain full repository awareness while operating within reasonable computational bounds—a critical requirement for production deployment.
III-E Deep Context Retrieval and Multi-Code Association
Chronos utilizes a novel multi-code compositional retrieval mechanism, powered by AGR, purpose-built for software engineering. Upon each reasoning or generation request—such as ”fix a failing test,” ”refactor authentication logic,” or ”explain API drift”—the Adaptive Retrieval Engine dynamically assembles a tailored context window by:
  • Issuing semantic queries to the Memory Engine that leverage both metric similarity and structural navigation in the code graph, with dynamic depth expansion (k-hop) based on query complexity.
  • Associating multiple code artifacts through typed relationships: e.g., tracing variable definitions across documentation (k=1), implementation (k=2), regression tests (k=2), and historic bug reports (k=3), stopping when confidence exceeds 90% or diminishing returns detected.
  • Dynamically refining the context through intermediate model inferences and confidence scoring, adapting retrieval depth in real-time. Complex debugging queries automatically trigger deeper graph traversal (k=3-5), while simple lookups terminate at k=1-2.
  • Utilizing edge type priorities: implementation edges (weight=1.0), dependency edges (weight=0.8), documentation edges (weight=0.6), ensuring most relevant paths are explored first.
TABLE II: Example multi-code association retrieval: constructing a task-specific context window for a bug fix.
Step Retrieved Entity Relationship
Q1 login.py Direct bug context
Q2 test_login.py Linked test
Q3 settings.py Imported env vars
Q4 bug_1234.md Historical bug doc
Q5 commit_a1b2c3 Last related commit
This approach allows Chronos to reason across arbitrarily distant, compositionally linked code and documentation artifacts—precisely what is needed for complex debugging, cross-module dependencies, or audit trails.
III-F Adaptive Graph-Guided Retrieval (AGR)
Traditional flat retrieval approaches fail to capture the intricate relationships between code artifacts, leading to incomplete context and erroneous fixes. Chronos introduces Adaptive Graph-Guided Retrieval (AGR), a dynamic mechanism that intelligently expands retrieval neighborhoods based on query complexity and confidence thresholds.
III-F1 Iterative Context Expansion
The AGR mechanism operates through iterative k-hop neighbor expansion:
  1. 1.
    Initial Query Analysis: Decompose the debugging request into semantic components and identify seed nodes in the code graph
  2. 2.
    Adaptive Depth Determination: Calculate optimal retrieval depth based on:
    • Query complexity score (0-1)
    • Code artifact density in the neighborhood
    • Historical debugging patterns for similar issues
  3. 3.
    Guided Expansion: Follow typed edges (implementation, dependency, dataflow) to retrieve contextually relevant nodes
  4. 4.
    Confidence-Based Termination: Stop expansion when retrieval confidence exceeds threshold or diminishing returns detected
III-F2 Graph-Guided vs Traditional Planning
Our empirical analysis reveals fundamental differences between traditional LLM planning and AGR-enhanced debugging:
[…]
Figure 4: Traditional LLM planning vs AGR-enhanced debugging: Graph-guided retrieval provides complete context, leading to accurate implementations.
k=1: Initial RetrievalPlanSignalImplTransEx1Ex2Retrieved:S1_next: Output signal(Type: Signal) k=2: Expanded RetrievalPlanSignalImplTransEx1Ex2StateWaitDocTestAdditional Retrieved:State transitions: Wait-¿S1Example: 9’b101000100100Confidence: 92%
Legend: Query k=1 k=2 Not Retrieved
Figure 5: Iterative context expansion in Adaptive Graph-Guided Retrieval: Starting from a query node, the system progressively expands retrieval depth (k-hops) based on confidence thresholds and query complexity.
BugVectorASTGraphHistoryRetrieverRankerSourceTestsFixesLogs
Figure 6: Multi-modal retrieval mechanism in Chronos.
III-G Reasoning, Generation, and Autonomous Orchestration
The transformer-based Chronos Reasoning Model operates directly over the retrieved, multi-source debugging context. Unlike classical code completion models, Chronos:
  • Diagnoses root causes and synthesizes code changes conditioned on project documentation, prior commits, and dependency patterns.
  • Produces stepwise fix plans, code diffs, documentation updates, and regression test suggestions in a unified, automated debugging loop.
  • Orchestrates a full debugging workflow: proposes bug fixes, invokes relevant tests, parses results, iterates on failures, and generates changelogs or PR summaries—all autonomously.
All outputs and feedback streams (test results, reviewer comments, CI/CD events) are fed back into the Memory Engine, enabling lifelong refinement and rapid adaptation to new debugging scenarios.
This cyclic process of context assembly, reasoning, autonomous validation, and memory update is the core of Chronos’s persistent codebase intelligence, enabling self-sustaining and ever-improving debugging at scale.
Autonomous Code/Documentation Generation(Bug Fix Proposal)AutomatedTesting & ValidationMemory UpdateIterative PlanRefinementTest PassedTest FailedReviewer FeedbackCI/CD Events
Figure 7: Chronos debugging feedback loop: Automated bug fix generation, validation, plan refinement, and memory update for continuous autonomous improvement.
IV THE DEBUGGING LANGUAGE MODEL PARADIGM
Kodezi Chronos fundamentally departs from traditional code models by being purpose-built as a debugging language model—the first of its kind. While existing LLMs treat debugging as a code generation problem, Chronos recognizes it as a complex, iterative process requiring specialized capabilities, training, and architecture.
IV-A Core Architectural Layers
Chronos implements a revolutionary 7-layer architecture specifically designed for autonomous debugging:
  1. 1.
    Multi-Source Input Layer: Ingests heterogeneous debugging signals including source code, CI/CD logs, error traces, stack dumps, configuration files, historical PRs, and issue reports. Unlike code models that primarily process source files, Chronos natively understands debugging artifacts.
  2. 2.
    Adaptive Retrieval Engine: Employs AGR (Adaptive Graph-Guided Retrieval) with a hybrid vector-symbolic approach combining:
    • Dynamic k-hop neighbor expansion based on query complexity
    • AST-aware code embeddings that preserve structural relationships
    • Dependency graph indexing for cross-file impact analysis
    • Call hierarchy mapping for execution flow understanding
    • Temporal indexing of code evolution and bug history
    • Confidence-based termination for optimal context assembly
  3. 3.
    Debug-Tuned LLM Core: A transformer architecture specifically fine-tuned on debugging workflows, not just code completion. Training tasks include:
    • Root cause prediction from symptoms
    • Multi-file patch generation
    • Test failure interpretation
    • Regression risk assessment
  4. 4.
    Orchestration Controller: Implements the autonomous debugging loop:
    • Hypothesis generation from error signals
    • Iterative fix refinement based on test results
    • Rollback mechanisms for failed attempts
    • Confidence scoring for proposed solutions
  5. 5.
    Persistent Debug Memory: Maintains long-term knowledge including:
    • Repository-specific bug patterns and fixes
    • Team coding conventions and preferences
    • Historical fix effectiveness metrics
    • Module-level vulnerability profiles
  6. 6.
    Execution Sandbox: Real-time validation environment supporting:
    • Isolated test execution
    • CI/CD pipeline emulation
    • Performance regression detection
    • Security vulnerability scanning
  7. 7.
    Explainability Layer: Generates human-readable outputs:
    • Root cause explanations with evidence chains
    • Fix rationale documentation
    • Automated PR descriptions and commit messages
    • Risk assessment reports
IV-B Training on Debugging Workflows
Unlike models trained primarily on code completion, Chronos’s training regime focuses exclusively on debugging scenarios:
Pre-training Corpus:
  • 15M+ GitHub issues with linked PRs and fix commits
  • 8M+ stack traces paired with resolutions
  • 3M+ CI/CD logs from failed and fixed builds
  • Production debugging sessions from enterprise partners
  • Open-source bug databases (Defects4J, SWE-bench, BugsInPy)
Specialized Fine-tuning Tasks:
  • Chain-of-Cause Reasoning: Teaching the model to trace error propagation through call stacks and dependencies [49]
  • Multi-Modal Bug Understanding: Correlating code, logs, traces, and documentation
  • Iterative Fix Refinement: Learning from failed fix attempts to improve subsequent proposals [50, 48]
  • Cross-Repository Pattern Recognition: Identifying similar bugs across different codebases
IV-C The Autonomous Debugging Loop
Chronos’s debugging loop represents a fundamental innovation over single-shot code generation:
Detect IssueRetrieve ContextPropose FixRun TestsTests Pass?Commit & DeployRefine StrategyUpdate MemoryYesNo
Figure 8: The Chronos autonomous debugging loop: continuous iteration until validation succeeds.
This loop continues autonomously, with each iteration informed by previous attempts and accumulated knowledge, until a validated fix is achieved or human intervention is requested.
V EVALUATION METHODOLOGY
To rigorously assess Kodezi Chronos’s capabilities across realistic debugging and maintenance workflows, we adopt a multi-faceted evaluation strategy that goes beyond conventional sequence completion or shallow retrieval tests.
V-A Benchmarking Datasets and Tasks
Chronos is evaluated on a diverse suite of benchmarks, comprising:
  • Standard Code Generation Benchmarks: HumanEval [2], MBPP [3], and related tasks for basic code synthesis and function-level reasoning.
  • Repository-Scale Debugging Tasks: A curated set of real-world bug reports, failing test scenarios, and codebase refactoring challenges drawn from open-source projects and industry case studies.
  • Multi-Code Association Retrieval: Custom synthetic benchmarks requiring retrieval and reasoning across arbitrarily placed, subtly linked code/documentation artifacts—simulating variable tracing, cross-file bug localization, or dependency update chains.
V-B Multi-Code Reasoning Evaluation Protocol
Unlike traditional benchmarks that target token-level prediction in narrow context, our protocol explicitly:
  • Randomizes the placement of relevant context (bug source, documentation clue, test assertion) across large codebases and histories.
  • Requires Chronos to retrieve, associate, and utilize multi-code context in a compositional manner, solving tasks that demand reasoning over both explicit code relationships (e.g., function calls, imports) and implicit bug/error propagation patterns.
  • Measures both retrieval accuracy (whether Chronos finds all necessary context) and end-to-end task success (whether it can autonomously fix, validate, and document the issue).
V-C Multi Random Retrieval Benchmark
We introduce the Multi Random Retrieval (MRR) benchmark, specifically designed to evaluate debugging-oriented retrieval capabilities:
V-C1 Benchmark Design
The MRR benchmark consists of 5,000 real-world debugging scenarios where:
  1. 1.
    Context Scattering: Relevant debugging information is randomly distributed across 10-50 files
  2. 2.
    Temporal Dispersion: Critical bug context spans 3-12 months of commit history
  3. 3.
    Obfuscated Dependencies: Variable names and function calls are refactored between bug introduction and discovery
  4. 4.
    Multi-Modal Artifacts: Solutions require combining code, tests, logs, and documentation
V-C2 Evaluation Metrics
  • Retrieval Precision@k: Fraction of retrieved artifacts that are relevant to the bug fix
  • Retrieval Recall@k: Fraction of all relevant artifacts successfully retrieved
  • Fix Accuracy: Whether the generated fix passes all tests and doesn’t introduce regressions
  • Context Efficiency: Ratio of used vs retrieved tokens in the final solution
V-C3 Results on MRR Benchmark
TABLE III: Performance on Multi Random Retrieval benchmark, demonstrating Chronos’s superior ability to find and utilize scattered debugging context.
Model Precision@10 Recall@10 Fix Accuracy Context Eff.
GPT-4 + RAG 42.3% 31.7% 8.9% 0.23
Claude-3 + Vector DB 48.1% 36.2% 11.2% 0.28
Gemini-1.5 + Graph 51.7% 41.8% 14.6% 0.31
Kodezi Chronos 89.2% 84.7% 67.3% 0.71
V-D Adaptive Graph-Guided Retrieval Performance
We evaluate the impact of AGR on debugging accuracy across different retrieval depths:
TABLE IV: Performance metrics for different retrieval strategies. Adaptive AGR dynamically selects optimal k based on query complexity.
Retrieval Method k=1 k=2 k=3 k=adaptive Flat
Precision 84.3±2.1% 91.2±1.4% 88.7±1.8% 92.8±1.2% 71.4±3.2%
Recall 72.1±2.8% 86.4±1.9% 89.2±1.6% 90.3±1.5% 68.2±3.5%
F1 Score 77.7±2.4% 88.7±1.6% 88.9±1.7% 91.5±1.3% 69.8±3.3%
Debug Success 58.2±3.1% 72.4±2.3% 71.8±2.4% 87.1±1.8% 23.4±4.1%
Key findings from AGR evaluation:
  • Optimal Depth Varies: Simple bugs require k=1-2, while complex cross-module issues benefit from k=3+
  • Adaptive Superiority: Dynamic depth selection outperforms fixed k values by 15-20%
  • 5x Improvement: AGR achieves 87.1% debug success vs 23.4% for flat retrieval
  • Hardware Debugging: Particularly effective for Verilog/VHDL with 91% accuracy (vs 18% baseline)
Model HumanEval MBPP Debug Success Root Cause Acc. Retrieval Prec.
GPT-4 85.7±1.2% 87.0±0.9% 8.5±2.1% 12.3±1.8% 68±2.3%
GPT-4-Turbo 87.1±1.0% 88.2±0.8% 9.2±1.9% 14.1±1.6% 71±2.0%
Claude-3-Opus 84.9±1.3% 86.1±1.1% 7.8±2.3% 11.7±2.0% 67±2.4%
Claude-3-Sonnet 83.2±1.4% 85.2±1.2% 7.3±2.4% 10.9±2.1% 66±2.5%
Gemini-1.5-Pro 88.3±0.9% 89.1±0.7% 11.2±1.7% 15.8±1.5% 74±1.8%
CodeT5+ 86.5±1.1% 84.8±1.3% 10.6±1.8% 13.2±1.7% 72±1.9%
Chronos 90.2±0.6%*** 88.9±0.5%* 65.3±1.4%*** 78.4±1.2%*** 91±0.8%***
*p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001 compared to best baseline (two-tailed t-test)
TABLE V: Performance across code synthesis and debugging tasks (mean ± std over 5 runs). Statistical significance shown.
V-E Comparison with Agentic Code Tools
While traditional LLMs struggle with debugging, a new generation of agentic code tools has emerged. We evaluate Chronos against these systems on real-world debugging scenarios:
Tool Context Memory Debug Loop Multi-File CI/CD Success Rate
Cursor Auto IDE only None No Limited No 4.2%
Claude Code 200K tokens Session No Yes No 6.8%
OpenAI Codex 8K tokens None No No No 3.1%
Gemini CLI 1M tokens None No Yes Limited 9.7%
GitHub Copilot X 16K tokens None No Limited No 5.3%
Chronos Unlimited* Persistent Yes Yes Yes 65.3%
TABLE VI: Comparison of Chronos with agentic code tools. *Unlimited via intelligent retrieval and memory.
Key differentiators of Chronos:
  • Persistent Memory: Unlike session-based tools, Chronos maintains cross-session knowledge of bugs, fixes, and patterns
  • True Debugging Loop: Automated iteration through fix-test-refine cycles until validation succeeds
  • CI/CD Integration: Native understanding of build systems, test frameworks, and deployment pipelines
  • Unlimited Context: Smart retrieval enables reasoning over entire repositories without token limits
V-F Cycle-Aware Debugging Case Study
To showcase Chronos’s real-world debugging prowess, we conduct a qualitative study involving a set of regression bug scenarios drawn from a large open-source Python project. Metrics include:
  • Number of attempts to converge on a passing code/test cycle.
  • Ability to document and explain root causes compared to human reviewers.
  • Time-to-resolution and reduction in manual engineering effort.
GPT-4ClaudeCodeT5+Chronos0022224444666688886.56.56.56.56.86.86.86.85.15.15.15.12.22.22.22.2ModelAvg. Cycles to Fix
Figure 9: Average code-to-fix cycles for Chronos and baseline models on real-world bugs (lower is better).
Chronos demonstrates not only higher accuracy and retrieval precision, but also a dramatically reduced number of debug cycles—underscoring its benefit for continuous, automated codebase reliability.
V-G Case Studies
To illustrate Chronos’s debugging capabilities, we present two real-world examples from our evaluation:
V-G1 Case Study 1: Cross-Module Null Pointer Exception
Bug Report: ”Application crashes with NullPointerException when processing user exports after recent authentication refactor”
Chronos Reasoning Process:
  1. 1.
    Retrieved recent commits mentioning ”authentication” (found 3 relevant commits)
  2. 2.
    Traced data flow from auth module to export module
  3. 3.
    Identified missing null check after auth token refresh
  4. 4.
    Located similar pattern in 2 other modules
Generated Fix:
Null Pointer Fix in Authentication Flow // In AuthService.java public AuthToken refreshToken(String userId) { AuthToken token = tokenCache.get(userId); if (token == null $|$ token.isExpired()) { token = authProvider.generateToken(userId); + if (token != null) { + tokenCache.put(userId, token); + } } return token; } // In ExportService.java public void exportUserData(String userId) { AuthToken token = authService.refreshToken(userId); + if (token == null) { + throw new AuthenticationException( + "Failed to authenticate user for export"); + } // ... rest of export logic }
Validation Results: All 47 test cases passed, including 3 new tests Chronos generated for the edge case.
V-G2 Case Study 2: Async Race Condition in Message Queue
Bug Report: ”Intermittent message loss in high-load scenarios, approximately 0.1% messages not processed”
Chronos Analysis:
  • Retrieved message queue implementation and recent load test results
  • Identified race condition between acknowledgment and connection pooling
  • Found similar issue fixed 8 months ago in different component
  • Applied analogous solution with adaptations
Generated Fix:
Race Condition Fix in Message Queue // MessageProcessor.java private void processMessage(Message msg) { try { handler.process(msg); - connectionPool.returnConnection(conn); - msg.acknowledge(); + // Fix: Acknowledge before returning connection + msg.acknowledge(); + connectionPool.returnConnection(conn); } catch (Exception e) { + // Ensure connection returned even on error + connectionPool.returnConnection(conn); msg.nack(); throw e; } }
Results: Load test with 10M messages showed 0
V-H Long-Context Debugging Performance
Even models with extended context windows fail at real debugging tasks due to fundamental architectural limitations:
Model Context Size Cross-File Bugs Historical Bugs Complex Traces Avg. Success
GPT-4-32K 32K tokens 7.2±1.9% 3.1±1.2% 5.8±1.6% 5.4±1.5%
Claude-3-200K 200K tokens 9.8±2.1% 4.7±1.4% 8.3±1.8% 7.6±1.7%
Gemini-1.5-Pro-1M 1M tokens 14.3±2.3% 6.2±1.5% 11.7±2.0% 10.7±1.9%
Chronos Unlimited 71.2±1.8%*** 68.9±2.0%*** 74.3±1.6%*** 71.5±1.8%***
***p ¡ 0.001 compared to Gemini-1.5-Pro-1M (paired t-test, n=100 tasks per category)
TABLE VII: Performance on debugging tasks requiring extensive context. Statistical significance demonstrates retrieval superiority.
The results demonstrate that raw context size alone cannot solve debugging. Chronos’s intelligent retrieval, persistent memory, and debug-specific training enable it to outperform even million-token models by over 6x.
V-I Detailed Performance Analysis
We further analyze Chronos’s performance across different bug categories and complexity levels:
Bug Category Syntax Logic Concurrency Memory API Performance
GPT-4 82.3% 12.1% 3.2% 5.7% 18.9% 7.4%
Claude-3-Opus 79.8% 10.7% 2.8% 4.3% 16.2% 6.1%
Gemini-1.5-Pro 85.1% 15.3% 4.1% 6.9% 22.4% 9.8%
Chronos 94.2% 72.8% 58.3% 61.7% 79.1% 65.4%
TABLE VIII: Success rates by bug category. Chronos shows particular strength in complex bug types that require deep understanding.
Repository Size ¡10K LOC 10K-100K 100K-1M ¿1M LOC
GPT-4 15.2% 9.8% 4.3% 1.2%
Claude-3-200K 17.8% 11.2% 5.7% 2.1%
Gemini-1.5-Pro 21.3% 14.7% 8.9% 3.8%
Chronos 71.2% 68.9% 64.3% 59.7%
TABLE IX: Debugging success rates by repository size, demonstrating Chronos’s scalability.
V-J Multi-Code Association Retrieval Performance
We evaluate Chronos’s ability to retrieve and associate multiple code artifacts for debugging:
Retrieval Task Precision Recall F1 Score
Variable Tracing 92.3±1.4% 89.7±1.6% 91.0±1.2%
Cross-File Dependencies 88.9±1.8% 91.2±1.5% 90.0±1.4%
Historical Bug Patterns 94.1±1.1% 87.3±2.0% 90.6±1.3%
Test-Code Mapping 91.7±1.3% 93.5±1.2% 92.6±1.0%
Documentation Links 85.4±2.1% 88.9±1.9% 87.1±1.7%
Average 90.5±0.8% 90.1±0.9% 90.3±0.7%
TABLE X: Multi-code association retrieval performance across different debugging contexts.
V-K Computational Efficiency
A critical consideration for production deployment is computational efficiency. We analyze Chronos’s performance characteristics compared to baselines and human debugging:
Metric GPT-4 Claude-3 Gemini-1.5 Chronos Human Dev
Avg. Time to Fix 82.3s 76.9s 71.2s 134.7s 2.4 hours
Context Window 128K tokens 200K tokens 1M tokens Unlimited* N/A
Cost per Bug \$0.47 \$0.52 \$0.68 \$0.89 \$180
Success Rate 8.5% 7.8% 11.2% 65.3% 94.2%
Effective Cost* \$5.53 \$6.67 \$6.07 \$1.36 \$191
*Unlimited via dynamic retrieval; Effective cost = Cost per bug / Success rate
TABLE XI: Computational efficiency and cost analysis. Despite higher per-attempt cost, Chronos’s high success rate yields lowest effective cost.
V-K1 Inference Time Breakdown
Chronos’s 134.7s average debugging time consists of:
  • Context Retrieval: 23.4s (17.4%)
  • Multi-round Reasoning: 67.8s (50.3%)
  • Test Execution: 31.2s (23.2%)
  • Memory Update: 12.3s (9.1%)
V-K2 Return on Investment Analysis
For a typical enterprise with 100 developers:
  • Annual debugging time:  150,000 hours
  • Chronos automation potential: 65.3% × 150,000 = 97,950 hours
  • Cost savings: 97,950 × \$90/hour - deployment costs = \$8.1M annually
  • ROI: 47:1 in first year, accounting for infrastructure and licensing
VI DISCUSSION AND ABLATION ANALYSIS
VI-A Qualitative Insights
Our evaluation highlights several domains where Chronos delivers outsized impact compared to prior systems:
  • Holistic Bug Localization: Chronos traces complex error origins across modules, commits, and documentation with no manual guidance, routinely identifying root causes overlooked by token-limited models.
  • Autonomous Debugging Loops: Chronos adapts its retrieval and patching behavior over multiple test cycles, integrating failed test feedback and reviewer commentary to iteratively refine solutions.
  • Continuous Knowledge Incorporation: By feeding CI/CD, reviewer, and test feedback into persistent memory, Chronos improves its project-specific performance over time, exhibiting lower repeated error rates and faster adaptation to new code patterns.
Bug Scenario GPT-4 Chronos Chronos Resolution Path
Test Failure on user_auth Incorrect var patch Full fix Traced import drift $\rightarrow$ found stale config $\rightarrow$ auto-fix and doc update
API Deprecation Missed call-site Full fix Multi-code association retrieved usage in 3 files, migrated all refs
Intermittent CI Error Flaky retry logic Full fix Ingested CI logs, patched async boundary, added test case and explanation
TABLE XII: Qualitative examples where Chronos successfully applies multi-code context to resolve debugging tasks beyond the reach of baseline LLMs.
VI-B Ablation Studies
To isolate the contribution of core design features, we perform targeted ablations:
  • No Multi-Code Association: When Chronos is restricted to single-chunk retrieval, debug success falls by 45% and retrieval precision drops sharply, mirroring the limitations of prior RAG pipelines.
  • Static Memory Only: If the live feedback/memory update mechanism is ablated (i.e., only static embeddings used), adaptivity stagnates, and repeated bug classes recur more often.
  • No Orchestration Loop: Disabling the validate-retrieve-update workflow reverts performance to basic code suggestion with higher error rate and longer time-to-fix.
FullNo MCAStaticNo Orch002020202040404040606060608080808010010010010090909090494949496262626255555555ConfigurationDebug Success (%)
Figure 10: Ablation analysis: Debugging success rate with each Chronos core component removed (lower is worse).
These findings underscore the essential synergy between deep memory, multi-code contextualization, and autonomous workflow orchestration for effective debugging and adaptive code maintenance.
VI-C Failure Analysis
Despite Chronos’s strong performance, our analysis reveals specific failure modes and bug categories where the system struggles:
VI-C1 Common Failure Modes
  1. 1.
    Hardware-Dependent Bugs: Chronos achieves only 23.4% success on bugs requiring hardware-specific knowledge (e.g., GPU memory alignment, embedded system timing). Example failure:
    Bug: CUDA kernel crashes with unaligned memory access on Tesla V100
    Chronos Fix: Added boundary checks (incorrect)
    Correct Fix: Aligned memory allocation to 128-byte boundaries
  2. 2.
    Distributed System Race Conditions: Complex timing-dependent bugs across multiple services show 31.2% success rate. The model struggles to reason about non-deterministic execution orders across network boundaries.
  3. 3.
    Domain-Specific Logic Errors: Bugs requiring deep domain knowledge (medical, financial regulations) succeed only 28.7% of the time. Example:
    Bug: HIPAA compliance violation in patient data export
    Issue: Chronos lacks healthcare regulatory knowledge
VI-C2 Edge Cases and Limitations
  • Extremely Large Monorepos (¿10M LOC): Performance degrades to 45.3% success rate due to retrieval precision issues
  • Legacy Code with Poor Documentation: Success drops to 38.9% when code lacks comments and uses cryptic variable names
  • Multi-Language Polyglot Systems: Cross-language bugs (e.g., Python calling Rust via FFI) show only 41.2% success
  • UI/UX Bugs: Visual rendering issues essentially unsolvable (8.3% success) without screenshot analysis
Bug Category Success Rate Primary Failure Reason
Hardware-Specific 23.4±3.2% Lacks hardware specs
Distributed Race 31.2±2.8% Non-deterministic timing
Domain Logic 28.7±3.1% Missing domain knowledge
Legacy Code 38.9±2.9% Poor code quality
Cross-Language 41.2±2.7% FFI complexity
UI/Visual 8.3±1.9% No visual understanding
TABLE XIII: Chronos performance on challenging bug categories.
VII LIMITATIONS, FUTURE WORK, AND BROADER IMPACT
VII-A Current Limitations
While Chronos marks a significant advance in autonomous code maintenance, certain limitations remain:
  • Extreme-Scale Context Latency: Despite highly efficient memory and retrieval, performance can be constrained in ultra-large monolithic repositories during peak concurrent update hours.
  • Initial Memory Cold Start: Brand new projects (no history) or projects with very sparse documentation may exhibit reduced multi-code association richness, necessitating a brief period of active learning.
  • Non-Determinism and Overcorrection: In some edge cases, especially with noisy historical feedback, Chronos has shown a tendency to ”over-correct” or introduce minor regressions before self-stabilization.
  • Opaque Reasoning Pathways: The fully autonomous nature and integration of deep learning sometimes hinder model explainability, which may limit adoption in highly regulated or safety-critical industries.
VII-B Future Work
We are actively addressing these limitations and have several concrete directions for improvement:
  • Incremental Embedding Optimization: Research into hybrid memory systems and smarter incremental re-indexing to further minimize retrieval and context assembly latency during scale-out operations.
  • Interactive and Visual Explanations: Implementation of stepwise reasoning visualizers and debug trace replays, making Chronos’ decision process more transparent to engineers and auditors.
  • Multi-Project/Organization Federated Memory: Extending Chronos memory to operate across multiple repositories and organizational boundaries, supporting federated bug detection, code migration, and systemic risk profiling.
  • Human-in-the-Loop Collaboration: Exploring adaptive workflows where engineers and Chronos can jointly control the loop—overriding, annotating, or confirming automated fixes in sensitive code regions.
  • Security and Adversarial Robustness: Building rigorous defense mechanisms against prompt injection, memory poisoning, and adversarial debugging cases.
VII-C Deployment Architecture and Integration
The proposed architecture extends beyond isolated debugging to comprehensive autonomous maintenance through a multi-tiered system design. The integration framework comprises:
  • Continuous Monitoring Layer: Real-time analysis of code quality metrics, security vulnerability patterns, and performance degradation indicators using static and dynamic analysis techniques
  • Automated Dependency Resolution: Graph-based impact analysis for dependency updates with probabilistic risk assessment and automated rollback mechanisms
  • Self-Healing Pipeline Integration: Event-driven architecture for autonomous incident response, incorporating validated patches into existing CI/CD workflows
  • Knowledge Synthesis Module: Automated extraction and formalization of implicit domain knowledge through documentation generation and code pattern analysis
Our empirical studies indicate that such integrated deployment can reduce mean time to resolution (MTTR) by 67% while maintaining a false positive rate below 3%. Field trials with industry partners are ongoing to validate these findings at scale.
VII-D Broader Impact
By enabling persistently self-healing and context-aware code maintenance, Chronos aims to shift an industry paradigm: reducing human toil and repetitive bug resolution, freeing engineers to focus on architecture, innovation, and user demands. As we scale deployment, it is crucial to steward responsible AI governance, data privacy, and an inclusive transition for developer workforces worldwide.
VIII CONCLUSION
We have presented Chronos, a novel debugging-specific language model that addresses fundamental limitations in existing code understanding systems. Through specialized training on debugging workflows and a purpose-built architecture incorporating persistent memory and intelligent retrieval, Chronos demonstrates significant improvements over general-purpose language models in automated debugging tasks.
Our comprehensive evaluation reveals that Chronos achieves a 65.3% success rate on real-world debugging benchmarks, representing a 6-7x improvement over state-of-the-art models including those with million-token contexts. This performance gain is attributed to three key technical contributions: (1) domain-specific pre-training on 15 million debugging instances including stack traces, fix commits, and CI/CD logs, (2) a persistent memory architecture that maintains cross-session knowledge of project-specific patterns, and (3) a hierarchical retrieval mechanism that enables effective reasoning over repository-scale contexts without computational constraints.
The implications of this work extend beyond immediate debugging applications. By demonstrating that specialized architectures and training regimes can dramatically improve performance on complex software engineering tasks, we provide evidence for the viability of task-specific language models in technical domains. Future research directions include extending this approach to other software engineering workflows, investigating transfer learning between debugging domains, and exploring human-AI collaborative debugging frameworks.
The transition toward autonomous debugging systems raises important considerations regarding software quality assurance, developer skill evolution, and the changing nature of software maintenance. As these systems mature, careful attention must be paid to maintaining human oversight, ensuring explainability of automated fixes, and preserving the creative and architectural aspects of software development that remain fundamentally human endeavors.
The Chronos model will be available in Q4 of 2025 and deploy on Kodezi [51] OS Q1 2026. This timeline allows for additional safety testing, enterprise integration development, and establishment of responsible deployment guidelines.
ACKNOWLEDGMENTS
Acknowledgments
This work benefited from the feedback and real-world challenges shared by early-access engineering partners, enterprise pilot users, and the broader Kodezi community. The maintainers of open-source repositories and tooling enabled large-scale benchmarking and inspired several retrieval and memory innovations described in this paper. Insights from researchers and practitioners in the software engineering and AI communities helped refine both methodology and experimental design. Support from Kodezi’s investors enabled the sustained research and development necessary to realize Chronos. Continued engagement and rigorous testing from the developer community have driven Chronos toward greater reliability and practical impact.

### v2

@@
 Abstract
-Large Language Models (LLMs) have advanced code generation and software automation, but are fundamentally constrained by limited inference-time context and lack of explicit code structure reasoning.
-We introduce Kodezi Chronos [51], a next-generation architecture for autonomous code understanding, debugging, and maintenance, designed to operate across ultra-long contexts comprising entire codebases, histories, and documentation—all without fixed window limits.
-Kodezi Chronos leverages a multi-level embedding memory engine, combining vector and graph-based indexing with continuous code-aware retrieval.
-This enables efficient and accurate reasoning over millions of lines of code, supporting repository-scale comprehension, multi-file refactoring, and real-time self-healing actions.
-Our evaluation introduces a novel Multi Random Retrieval benchmark, specifically tailored to the software engineering domain.
-Unlike classical retrieval benchmarks, this method requires the model to resolve arbitrarily distant and obfuscated associations across code artifacts, simulating realistic tasks such as variable tracing, dependency migration, and semantic bug localization.
-Chronos outperforms prior LLMs and code models—demonstrating a 23% improvement in real-world bug detection and reducing debugging cycles by up to 40% compared to traditional sequence-based approaches.
-By natively interfacing with IDEs and CI/CD workflows, Chronos enables seamless, autonomous software maintenance, elevating code reliability and productivity while reducing manual effort.
-These results mark a critical advance toward self-sustaining, continuously optimized software ecosystems.
+[…]
+On 5,000 real-world scenarios, Chronos achieves 67.3% ± 2.1% fix accuracy versus 14.2% ± 1.3% (Claude 4) and 13.8% ± 1.2% (GPT-4.1), with Cohen’s d=3.87 effect size.
+The system reduces debugging time by 40% and iterations by 65%.
+Chronos resolves complex multi-file bugs requiring cross-repository understanding and temporal analysis.
+Key limitations include 23.4% success on hardware-dependent bugs and 41.2% on dynamic language issues.
+Theoretical analysis proves O(k log d) retrieval complexity with convergence guarantees.
+Human evaluation (N=50) shows 89% preference over baselines.
+Available Q4 2025 (Kodezi OS) and Q1 2026 (API).
 I INTRODUCTION
 Recent advancements in large language models (LLMs) have transformed code generation, review, and reasoning tasks [1, 2, 3].
-However, debugging—the most time-consuming and critical aspect of software development—remains largely unsolved.
-While tools like GitHub Copilot [11], Cursor, and Claude excel at code completion [12], they fundamentally misunderstand debugging as a multi-faceted, context-heavy process that spans entire repositories, historical commits, CI/CD logs, and runtime behaviors.
+In 2025, frontier models like Claude Opus 4 [4] achieve 72.5% on SWE-bench, GPT-4.1 [5] reaches 54.6%, and DeepSeek V3 [6] demonstrates remarkable efficiency with only \$5.6M training cost.
+However, debugging, the most time-consuming and critical aspect of software development, remains largely unsolved.
+While tools like GitHub Copilot [7, 8], Cursor, Windsurf [9], and Claude excel at code completion [10], they fundamentally misunderstand debugging as a multifaceted, context-heavy process that spans entire repositories, historical commits, CI/CD logs, and runtime behaviors.
 Production debugging requires reasoning across files separated by thousands of lines [11], understanding temporal code evolution, and correlating seemingly unrelated symptoms to root causes buried deep in dependency chains [12].
-Current code assistants fail at debugging for three critical reasons: (1) they are trained primarily on code completion tasks, not debugging workflows [15]; (2) they lack persistent memory of past bugs, fixes, and codebase-specific patterns [16]; and (3) their context windows, even when extended to 100K+ tokens, cannot capture the full debugging context needed for complex, multi-file issues.
-Recent studies show that even state-of-the-art models like GPT-4, Claude-3, and Gemini-1.5 achieve less than 10% success rates on real-world debugging benchmarks [17, 18], often proposing superficial fixes that fail validation or introduce new regressions [19].
+Why does debugging remain fundamentally unsolved for LLMs?
+Beyond the surface-level challenges, debugging exposes four core limitations of current architectures.
+First, hallucination under uncertainty: when LLMs encounter ambiguous error states, they confidently generate plausible-sounding but incorrect fixes, often introducing subtle bugs that pass initial tests but fail in production.
+Second, absence of persistent memory: each debugging session starts tabula rasa, unable to learn from previous encounters with similar bugs or build institutional knowledge about codebase-specific patterns.
+Third, rigid token limits: even 1M-token contexts cannot capture the full transitive closure of dependencies in modern software, where a bug’s root cause may span dozens of files connected through deep call chains.
+Fourth, lack of causal reasoning: LLMs excel at pattern matching but struggle with counterfactual reasoning (”what if this variable were null?”) and temporal causality (”which commit introduced this regression?”), both essential for debugging.
+Current code assistants fail at debugging for three critical reasons: (1) they are trained primarily on code completion tasks, not debugging workflows [13]; (2) they lack persistent memory of past bugs, fixes, and codebase-specific patterns [14]; and (3) their context windows, even when extended to 1M tokens (Gemini 2.0) or leveraging advanced RAG techniques like HyDE [15] and FLARE [16], cannot capture the full debugging context needed for complex, multi-file issues.
+Recent studies show that even state-of-the-art models like GPT-4.1, Claude 4 Opus, and Gemini 2.0 Pro achieve less than 15% success rates on real-world debugging benchmarks (COAST, MLDebugging, MdEval) [17, 18], often proposing superficial fixes that fail validation or introduce new regressions [19].
 Kodezi Chronos represents a paradigm shift: the first debugging language model developed by Kodezi [20], specifically designed, trained, and optimized for autonomous bug detection, root cause analysis, and validated fix generation.
-Unlike code completion models that generate syntactically correct but often semantically flawed suggestions, Chronos operates through a continuous debugging loop—proposing fixes, running tests, analyzing failures, and iteratively refining solutions until validation succeeds.
-Built on a novel architecture combining persistent debug memory, multi-source retrieval (code, logs, traces, PRs), and execution sandboxing, Chronos achieves what no existing model can: true autonomous debugging at repository scale.
-Chronos is designed for seamless integration with modern development stacks: it operates as an embedded “AI CTO” within CI/CD pipelines, IDEs, and project management tools.
+For more information about the model and benchmarks, visit https://chronos.so/ and https://github.com/kodezi/chronos.
+Kodezi OS information is available at https://kodezi.com/os.
+Unlike code completion models that generate syntactically correct but often semantically flawed suggestions, Chronos operates through a continuous debugging loop, proposing fixes, running tests, analyzing failures, and iteratively refining solutions until validation succeeds.
+Built on a novel architecture combining persistent debug memory, multi-source retrieval (code, logs, traces, PRs), and execution sandboxing, Chronos demonstrates debugging performance that significantly exceeds current models, particularly in repository-scale, multi-file scenarios.
+Chronos is designed for seamless integration with modern development stacks: it operates as an embedded autonomous maintenance system within CI/CD pipelines, IDEs, and project management tools.
 Its proactive, event-driven workflow ensures that debugging, documentation generation, refactoring, and even preventative maintenance actions occur autonomously, guided by deep repository memory rather than manual prompts or brittle heuristics.
@@
 Empirically, Chronos demonstrates state-of-the-art results across industry-standard metrics and realistic maintenance tasks, reducing debugging times and increasing project resilience.
-This paper presents the architecture, memory system, retrieval mechanism, evaluation methodology, and experimental results that establish Kodezi Chronos as the new frontier for autonomous, ultra-contextual codebase management.
-II RELATED WORK AND LIMITATIONS OF TRADITIONAL APPROACHES
-The emergence of large-scale neural models for source code processing—such as CodeBERT [46], GraphCodeBERT [4], and CodeT5 [5]—has significantly advanced automatic code synthesis, translation, and review.
-Many of these models, as well as massive LLMs like GPT-3 and GPT-4 [1, 47], are pre-trained on billions of lines of code paired with natural language, learning rich semantic representations for many programming languages.
+This paper presents the architecture, memory system, retrieval mechanism, evaluation methodology, and experimental results that establish Kodezi Chronos as the new frontier for autonomous debugging with full repository context.
+Why Chronos Dominates 87.1% debugging success vs.
+22.9% for GPT-4.1 + SWE-agent 3.8x faster bug resolution through AGR’s precision retrieval 67% fewer false positives via Persistent Debug Memory (PDM) First to handle cross-file, temporal, and runtime-dependent bugs
+Our contributions are as follows:
+  1. 1.
+    We introduce Chronos, a debugging-specific language model architecture integrating persistent memory, adaptive graph-based retrieval (AGR), and a multi-stage fix orchestration loop.
+  2. 2.
+    We design AGR, a multi-hop, edge-weighted graph traversal system for deep context retrieval tailored to codebases, achieving 92% precision at 85% recall on debugging queries.
+  3. 3.
+    We construct new debugging-specific evaluation benchmarks, including cycle-aware test loops, retrieval accuracy, and root-cause localization across 12,500 real-world bugs.
+  4. 4.
+    We demonstrate significant performance gains over Claude 4 Opus, GPT-4.1, and specialized tools, achieving 65.3% debugging success rate on our comprehensive benchmarks (4-5x improvement over GPT-4.1, Claude 4 Opus).
+  5. 5.
+    We provide comprehensive ablation studies showing each component’s contribution, with detailed case studies and quantitative analysis demonstrating 4-5x improvement over state-of-the-art models.
+II RELATED WORK: LIMITATIONS OF EXISTING APPROACHES IN DEBUGGING
+II-A Retrieval-Augmented Code Generation
+The emergence of large-scale neural models for source code processing, such as CodeBERT [21], GraphCodeBERT [22], and CodeT5 [23], has significantly advanced automatic code synthesis, translation, and review.
+CodeT [24] introduced code generation with testing, improving reliability through execution feedback.
+Many of these models, as well as massive LLMs from the GPT-4 family [1, 5], are pre-trained on billions of lines of code paired with natural language, learning rich semantic representations for many programming languages.
+State-of-the-art retrieval-augmented generation (RAG) methods have evolved significantly in 2025.
+Advanced techniques like HyDE (Hypothetical Document Embeddings) [15] generate synthetic documents to improve retrieval quality, Self-RAG [25] uses reflection tokens for dynamic retrieval decisions, and FLARE [16] implements forward-looking active retrieval based on generation confidence.
+GraphRAG [26] integrates knowledge graphs for structured retrieval, representing a step toward graph-based understanding but still operating with static graphs that lack the dynamic updates required for debugging workflows.
+Recent work on code-specific retrieval [27] shows promise but lacks the multi-hop reasoning capabilities essential for debugging.
+LangChain [28] and DSPy [29] provide frameworks for building complex LLM applications with retrieval components.
+While these frameworks enable sophisticated pipelines, they fundamentally rely on stateless retrieval without the persistent memory or specialized debugging knowledge that Chronos provides.
+DSPy’s optimization of prompting strategies and LangChain’s chain-of-thought orchestration, while powerful for general tasks, lack the domain-specific understanding of code dependencies and bug patterns that debugging requires.
 Despite impressive gains on benchmark tasks, these approaches are fundamentally bottlenecked by attention-based architectures and fixed-size input windows, typically constraining context to tens of thousands of tokens.
-Practical developer workflows, by contrast, require tools to reason across entire repositories: files, modules, historic commits, documentation, and evolving build systems, far outstripping standard context limits.
-State-of-the-art retrieval-augmented generation (RAG) methods extend practical context by embedding and recalling external documents [8], but remain mostly limited to chunked passage retrieval and lack true compositional, dependency-aware memory.
-Window expansion techniques—such as models with 100k tokens or more [7]—incur prohibitive compute/memory costs and suffer from diluted attention, leading to information loss and degraded performance as codebase size increases.
-Existing benchmarks for long-context reasoning, like the ”Needle in a Haystack” pattern, commonly involve embedding a salient but easily differentiable clue in a sea of noise or repetition.
-Such evaluations unintentionally rely on the model’s ability to match explicit tokens or unusual patterns, providing only a partial test of true retrieval and reasoning ability.
-In real-world maintenance, however, code elements are frequently unmarked, distributed, and semantically interdependent, requiring the AI not just to ”find,” but to infer context, resolve relationships, and reason about impact [45].
+Window expansion techniques, such as models with 1M+ tokens (Gemini 2.5, GPT-4.1) [30, 5], incur prohibitive compute/memory costs and suffer from diluted attention, leading to information loss and degraded performance as codebase size increases.
+II-B Program Repair and Bug Localization
+Recent systems targeting debugging include SWE-agent [13], which achieves 12.3% on the SWE-bench dataset through structured agent-computer interfaces, and AutoCodeRover [17], reaching 22.9% via program structure-aware retrieval.
+These systems represent significant progress but still operate without persistent memory across debugging sessions.
+The SWE-bench family of benchmarks has expanded significantly: SWE-bench++ [31] introduces harder real-world scenarios with multi-repository dependencies, while SWE-bench-coding [32] focuses on implementation tasks rather than bug fixes.
+Despite these advances, even state-of-the-art models struggle with the full complexity of real-world debugging, achieving less than 25% success on these enhanced benchmarks.
 Other recent work explores graph neural networks (GNNs) for modeling explicit data flow or control flow graphs [33, 34].
 While GNNs can capture structural properties of code, they are rarely coupled with ultra-long context LLMs, and lack the continuous learning, memory updating, and rapid recall required for live autonomous maintenance.
+The debugging challenge is further highlighted by specialized benchmarks: DebugBench [35] evaluates root cause analysis, while BugHunter [36] tests multi-file bug localization.
+However, these benchmarks often simplify real-world debugging by providing isolated test cases rather than full repository contexts.
+II-C Multi-Agent Systems for Software Tasks
+The rise of multi-agent architectures has introduced new paradigms for complex software tasks.
+Systems like AutoGPT [37] and BabyAGI [38] demonstrate task decomposition and autonomous execution, but lack the specialized knowledge required for debugging.
+MetaGPT [39] simulates software company workflows with multiple specialized agents, yet still operates without persistent cross-session memory.
+LangGraph [40] enables building stateful, multi-actor applications with LLMs, providing infrastructure for complex agent interactions.
+When combined with ReAct [41] loops, these systems can perform iterative reasoning.
+However, our evaluation shows that even LangGraph + ReAct configurations achieve only 31.2% debugging success compared to Chronos’s 65.3%, primarily due to: 1.
+Lack of persistent memory: Each debugging session starts fresh without learning from past fixes 2.
+Generic reasoning: No specialized understanding of debugging patterns or code dependencies 3.
+Inefficient exploration: Without AGR’s guided traversal, agents waste cycles on irrelevant code paths
+ChatDev [42] and similar multi-agent coding systems excel at greenfield development but struggle with the complexity of debugging existing codebases.
+The key differentiator is that Chronos operates with continuous memory updates, specialized debugging knowledge, and efficient graph-guided retrieval, capabilities absent in generic multi-agent frameworks.
+II-D How Chronos Differs
+Chronos distinguishes itself from existing approaches through three fundamental architectural decisions:
+1.
+Memory Scale: While LangChain and similar frameworks operate with session-level memory, Chronos maintains persistent debug memory (PDM) across millions of debugging sessions, learning and adapting from each interaction.
+2.
+Debug Reasoning: Unlike generic multi-agent systems that apply general problem-solving strategies, Chronos is trained specifically on 15M+ debugging scenarios, understanding patterns like race conditions, memory leaks, and API migrations that generic models miss.
+3.
+Orchestration: Rather than relying on user-driven or generic agent loops, Chronos implements a specialized 7-layer debugging architecture with automatic test validation, iterative refinement, and confidence-based termination.
+The combination of these capabilities enables Chronos to achieve 4-5x better debugging performance than state-of-the-art alternatives, as demonstrated in our comprehensive evaluation.
+II-E Performance Analysis: Code Generation vs Debugging Tasks
+The landscape of code-focused LLMs has evolved dramatically in 2025.
+Claude Opus 4 and Sonnet 4 [4] achieve 72.5% and 72.7% respectively on SWE-bench, representing the current state-of-the-art for code generation tasks.
+GPT-4.1 [5] doubles GPT-4o’s performance on code diff benchmarks and reaches 54.6% on SWE-bench tasks.
+Gemini 2.5 Pro [30] achieves 63.8% on SWE-bench with a custom agent setup, showcasing Google’s advances in reasoning models.
+DeepSeek V3 [6], with 671B parameters (37B activated), demonstrates that efficient training is possible, achieving competitive performance at 1/10th the training cost of comparable models.
+Qwen2.5-Coder-32B [43] matches GPT-4o performance while running locally on consumer hardware.
+However, these impressive code generation capabilities do not translate to debugging success.
+When evaluated on real-world debugging tasks requiring multi-file understanding, historical context, and iterative refinement (using COAST, MLDebugging, and MdEval benchmarks), even the best models achieve less than 15% success rates.
+This gap between code generation and debugging performance motivates the need for specialized debugging-focused architectures.
+TABLE I: Comparison of debugging approaches across different systems and their key limitations
+Approach Context Memory Debug Training Iteration Graph Key Limitation
+General LLMs
+Claude 4/GPT-4.1 200K-1M Session $\times$ $\times$ $\times$ No debug specialization
+Gemini 2.0 Pro 2M Session $\times$ $\times$ $\times$ Attention dilution
+RAG Approaches
+HyDE/Self-RAG Unlimited* None $\times$ $\times$ $\times$ Static retrieval
+FLARE Unlimited* None $\times$ Limited $\times$ No debug signals
+Graph RAG Unlimited* Static $\times$ $\times$ Static graphs
+Code Systems
+SWE-agent Limited None Partial $\times$ No persistent memory
+AutoCodeRover Unlimited* None Partial $\times$ Structure-only retrieval
+Cursor/Windsurf Session Session $\times$ Manual $\times$ User-driven loops
+Specialized
+Chronos Unlimited* Persistent Auto -
+*Through intelligent retrieval, not raw context window
 Kodezi Chronos is motivated by these challenges: By combining continuous graph-aware indexing, dynamic embedding updates, and reasoning-optimized memory retrieval, Chronos transcends traditional limitations and enables truly repository-scale, real-time software comprehension and intervention.
-III ARCHITECTURE OF KODEZI CHRONOS
-III-A The Output-Heavy Nature of Debugging
+III CHRONOS ARCHITECTURE: DESIGN AND IMPLEMENTATION
+This section presents the technical architecture of Chronos, beginning with the fundamental insight that debugging is output-heavy rather than input-heavy, followed by the core architectural components and implementation details.
+III-A Debugging as an Output-Heavy Task
 Despite the industry focus on ever-larger context windows (128K, 200K, 1M+ tokens), debugging presents a fundamentally different challenge: it is inherently output-heavy rather than input-heavy.
@@
   • Prior fix attempts: 500-1K tokens
-  • Total input: Often ¡10K tokens for most real-world debugging tasks
+  • Total input: Often $<10K$ tokens for most real-world debugging tasks
 What models must produce (output):
@@
 III-A2 Why Output Quality Trumps Input Size
-TABLE I: Input vs output characteristics in debugging tasks.
+TABLE II: Input vs output characteristics in debugging tasks.
 Aspect Input Context Output Generation
@@
 The critical insight: a model with intelligent 8K context that generates robust, test-passing fixes will outperform a 1M-context model that produces syntactically correct but semantically flawed patches.
+Debugging Token Flow: Input vs OutputInput (Sparse)Stack Trace: 300 tokensCode Context: 2K tokensLogs/Tests: 800 tokensHistory: 500 tokensTotal: $\sim$3.6K tokensProcessOutput (Dense)Bug Fix: 1.2K tokensExplanation: 500 tokensTests: 600 tokensDocs/PR: 400 tokensFallback: 300 tokensTotal: $\sim$3K tokensKey Insight: Output $\approx$ Input in debuggingQuality matters more than quantity
+Figure 1: Token distribution in debugging tasks: Unlike typical LLM applications where input dominates, debugging requires substantial, high-quality output generation.
 III-A3 Chronos’s Output-Optimized Architecture
@@
   2. 2.
-    Iterative Refinement Loop: Rather than single-shot generation, Chronos validates outputs through execution, using test results to refine patches—ensuring output quality over quantity.
+    Iterative Refinement Loop: Rather than single-shot generation, Chronos validates outputs through execution, using test results to refine patches, ensuring output quality over quantity.
   3. 3.
@@
 This output-centric design enables Chronos to achieve 65.3% debugging success despite competitors having 10-100x larger context windows, validating that for debugging, output quality and structure matter more than input capacity.
-III-B Core Architecture Overview
+Context RetrievalMemory UsageTest IntegrationMulti-File SupportError AnalysisFix GenerationIteration SpeedCost Efficiency20%40%60%80%100%ChronosClaude 4 OpusGPT-4.1
+Figure 2: Debugging capability comparison across eight key factors.
+Chronos (green) significantly outperforms general-purpose models in all dimensions, with particularly strong advantages in test integration (92%), iteration speed (95%), and cost efficiency (91%).
+III-B High-Level Architecture: From Error Signal to Validated Fix
 Kodezi Chronos is designed as an autonomous memory-driven intelligence layer for code, operating at scales that span entire enterprise repositories, team histories, and auxiliary knowledge sources.
-Its architecture consists of three core modules: (i) a persistent Memory Engine for continuous graph-based context construction, (ii) an advanced Retriever that dynamically assembles deep semantic context from code and documentation, and (iii) a transformer-based Code Reasoning Model for synthesis, debugging, and orchestration of software changes.
+Its architecture consists of three core modules: (i) a persistent Memory Engine for continuous graph-based context construction, (ii) an advanced Retriever that constructs targeted context from code and documentation, and (iii) a transformer-based Code Reasoning Model for synthesis, debugging, and orchestration of software changes.
+YesNoBug TriggeredPDMAccessAGRRetrievalChronos PlanGenerationContextAssemblyPatchDraftedTest LoopExecutedPass?Commit+ Update PDMRefine+ RetryAvg 2.2iterationsMemorypersists
+Figure 3: Complete fix loop lifecycle showing integration between PDM, AGR retrieval, and iterative refinement.
+Dashed lines indicate feedback mechanisms that enable learning across debugging sessions.
+III-C The Four Architectural Pillars
+Chronos demonstrates improved debugging performance through four architectural components that distinguish it from general-purpose code models:
+III-C1 Debugging-Specific Training on 15M+ Real Sessions
+Unlike models trained on static code repositories, Chronos learns from:
+  • 15 million debugging sessions from production environments
+  • Complete fix trajectories: initial bug report $\rightarrow$ attempted fixes $\rightarrow$ test failures $\rightarrow$ successful resolution
+  • Failure patterns: Common anti-patterns, regression indicators, and fix validation strategies
+  • Domain-specific knowledge: Framework quirks, library-specific debugging techniques, language idioms
+This specialized training enables Chronos to recognize subtle bug patterns that general models miss.
+For example, when encountering a React hydration mismatch, Chronos immediately knows to check server/client rendering differences rather than pursuing surface-level fixes.
+III-C2 Execution Sandbox with Real-Time Feedback Loop
+Chronos operates within a sophisticated execution environment that provides:
+  • Isolated test execution: Every proposed fix runs in a containerized sandbox
+  • Comprehensive validation: Unit tests, integration tests, linting, type checking
+  • Iterative refinement: Failed fixes generate detailed error logs fed back into the next iteration
+  • Regression prevention: Automatic detection of new failures introduced by fixes
+This execution-driven approach explains why Chronos averages 7.8 iterations per bug while general models stop after 1-2 attempts.
+Each iteration refines understanding based on concrete execution results rather than probabilistic guessing.
+III-C3 Persistent Repository Memory Across Sessions
+Unlike stateless models that start fresh each time, Chronos maintains:
+  • Bug pattern database: Historical bugs, their root causes, and successful fixes
+  • Codebase evolution graph: How files, functions, and dependencies changed over time
+  • Team-specific patterns: Coding conventions, common mistakes, architectural decisions
+  • Dependency knowledge: Version-specific quirks, migration paths, compatibility issues
+When debugging a null pointer exception, Chronos recalls similar bugs from 6 months ago, checks if recent refactoring introduced the issue, and applies team-specific null-safety patterns.
+III-C4 Adaptive Graph-Guided Retrieval (AGR) for Multi-File Context
+AGR enables Chronos to navigate complex codebases through:
+  • Dynamic graph construction: Real-time building of dependency graphs during debugging
+  • Intelligent k-hop traversal: Adaptively expanding search radius based on bug complexity, leveraging graph attention mechanisms [44]
+  • Semantic + structural retrieval: Combining code semantics with architectural relationships
+  • Temporal awareness: Understanding when code changed relative to bug introduction
+This sophisticated retrieval explains Chronos’s 89.2% precision on the MRR benchmark, it finds the needle in the haystack by understanding the haystack’s structure.
+TABLE III: Impact of each architectural pillar on debugging performance. Ablation study shows cumulative benefits.
+Configuration MRR Fix Acc. SWE-bench COAST Avg. Iterations Time (min)
+Base Model (no pillars) 8.3% 12.1% 9.7% 1.2 8.5
++ Debug Training 24.7% 28.3% 26.4% 2.8 16.2
++ Execution Sandbox 41.2% 43.7% 42.8% 5.1 28.4
++ Persistent Memory 55.8% 57.2% 56.3% 6.4 35.7
++ AGR (Full Chronos) 67.3% 65.3% 67.5% 7.8 42.3
 Code, Docs, CI/CD LogsMemory Engine(Embedding + Graph)Multi-CodeAssociation RetrieverReasoning Model& OrchestrationTest ResultsPatches, Changelogs,Test Results
 Figure 4: High-level overview of Chronos: Memory-driven embedding and retrieval powering autonomous reasoning and codebase management.
-[…]
-Figure 2: Token distribution in debugging tasks: Unlike typical LLM applications where input dominates, debugging requires substantial, high-quality output generation.
-III-C Memory Engine: Repository-Scale Embedding and Indexing
-The Memory Engine is responsible for ingesting, encoding, and maintaining a unified semantic representation of all project files, versioned code, documentation, configuration, historical diffs, test outcomes, and architectural artifacts.
-Each unit of code (e.g., function level, file, commit) is parsed to extract both sequence and structural cues—using ASTs, dependency graphs, and metadata links.
-These elements are then projected into a high-dimensional vector space via context-aware, code-specific encoders trained on multiple programming languages and coding paradigms.
-The Memory Engine stores not just static embeddings but also maintains an evolving graph database, where nodes represent code elements and edges denote code relationships (e.g., function calls, module imports, bug-ticket links, commit ancestry).
+III-D Persistent Debug Memory (PDM): Learning from Historical Fixes
+The Persistent Debug Memory (PDM) provides the foundation for Chronos’s persistent, repository-scale understanding.
+Unlike traditional approaches that recompute context for each query, PDM maintains a continuously updated semantic representation of the entire codebase, bug patterns, and fix history.
+III-D1 Unified Semantic Representation
+PDM ingests and encodes diverse artifacts including source code, documentation, configuration files, historical diffs, test outcomes, and architectural specifications.
+Each code unit (function, class, module, commit) undergoes multi-level analysis:
+  • Syntactic parsing: AST extraction for structural understanding
+  • Semantic embedding: Context-aware vectorization using specialized encoders
+  • Relational mapping: Graph construction capturing dependencies, calls, and evolution
+  • Temporal indexing: Version-aware representation enabling historical analysis
+TABLE IV: Persistent Debug Memory (PDM) Architecture and Policies
+Component Details
+Data Storage
+Code Snapshots Full AST + semantic embeddings per commit
+Bug Patterns Failed fixes, error signatures, stack traces
+Fix History Successful patches with test validation results
+CI/CD Logs Build failures, test outputs, deployment issues
+Documentation README, comments, design docs, PRs
+Retention Policy
+Active Bugs Permanent until resolved + 90 days
+Successful Fixes Permanent (forms learning corpus)
+Code Versions Last 1000 commits or 2 years
+Test Results 180 days rolling window
+Embeddings Re-computed weekly, cached 30 days
+Retrieval Policy
+Primary Index Semantic similarity (cosine distance)
+Secondary Index Temporal proximity + file dependencies
+Ranking Bug recency × pattern frequency × fix success
+Context Window Adaptive 5-50 nodes based on confidence
+Update Triggers
+Git Events Commit, merge, rebase (real-time)
+CI/CD Events Test failure, build break (<1<1< 1 min)
+Bug Reports Issue creation/update (<5<5< 5 min)
+Fix Validation Successful test run (immediate)
+Scheduled Full re-indexing (weekly)
+III-D2 Graph-Based Knowledge Storage
+The Memory Engine maintains an evolving graph database where nodes represent code elements and edges denote relationships.
+This dual vector-graph representation enables both semantic similarity search and structural traversal, providing the foundation for multi-hop reasoning during debugging.
+III-D3 PDM Retrieval Mechanism
+The Persistent Debug Memory employs a hybrid retrieval strategy combining temporal, semantic, and structural signals:
+  • Temporal-Aware Retrieval: Recent bugs and fixes are weighted higher (decay factor: $e^{-\lambda t}$, $\lambda=0.1$)
+  • Semantic Vector Search: FAISS index with 768-dim embeddings, cosine similarity threshold 0.75
+  • Graph Traversal: BFS/DFS from error location with typed edge filtering (imports, calls, inherits)
+  • Pattern Matching: Regex-based search for error signatures, stack trace patterns
+  • Hybrid Scoring: $Score=0.4\cdot Semantic+0.3\cdot Temporal+0.2\cdot Structural+0.1\cdot Pattern$
+This multi-modal retrieval ensures PDM surfaces relevant debugging context even when exact matches don’t exist, enabling cross-bug learning and pattern recognition.
 login.pyauth_utils.pyREADME.mdtest_login.pybug_1234.mdcallscited inlinked bug intested by
@@
 This design enables Chronos to efficiently retrieve, traverse, and reason about segments of the codebase that share non-local relationships, even if separated by thousands of lines, multiple files, or extensive revision history.
-III-D Achieving Unlimited Context Through Smart Retrieval
+III-E Breaking Token Limits: Intelligent Retrieval at Repository Scale
 Traditional LLMs are fundamentally constrained by attention complexity and memory limitations.
@@
 Chronos implements true unlimited context through:
-  • Hierarchical Code Embeddings: Multi-level representations from token → statement → function → module → repository
+  • Hierarchical Code Embeddings: Multi-level representations from token $\rightarrow$ statement $\rightarrow$ function $\rightarrow$ module $\rightarrow$ repository
   • Temporal Context Indexing: Every code element tagged with commit history, allowing time-travel debugging
@@
   • Dynamic Context Assembly: At inference, retrieves precisely the code paths relevant to the current bug
-This approach enables Chronos to maintain full repository awareness while operating within reasonable computational bounds—a critical requirement for production deployment.
-III-E Deep Context Retrieval and Multi-Code Association
-Chronos utilizes a novel multi-code compositional retrieval mechanism, powered by AGR, purpose-built for software engineering.
-Upon each reasoning or generation request—such as ”fix a failing test,” ”refactor authentication logic,” or ”explain API drift”—the Adaptive Retrieval Engine dynamically assembles a tailored context window by:
+This approach enables Chronos to maintain full repository awareness while operating within reasonable computational bounds, a critical requirement for production deployment.
+III-E1 How Chronos Achieves AGR Cost-Effectively
+While implementing AGR-style retrieval might seem computationally prohibitive, Chronos achieves it efficiently through five key architectural decisions:
+1. Debugging-Only Focus: Unlike general-purpose assistants handling millions of arbitrary queries, Chronos optimizes exclusively for fix loops. This narrow scope means:
+  • Fewer retrievals per session (avg 3.7 vs 15+ for autocomplete)
+  • Higher value per retrieval (each must trace root cause)
+  • Better cost-benefit ratio per inference
+2. One-Time Graph Construction: Chronos builds the repository graph once and updates incrementally:
+  • Initial AST+dependency parsing: 2-4 hours per 1M LOC
+  • Incremental updates on commits: <100<100< 100ms per file
+  • Graph reused across all debugging sessions
+  • Cost amortizes efficiently over many debugging sessions
+3. Smart Caching and Memory:
+  • PDM caches successful traversal paths (87% hit rate on recurring bugs)
+  • Common subgraphs pre-computed and indexed
+  • Frequently accessed nodes kept in hot storage
+  • Result: 47ms retrieval for cached patterns vs 3.2min cold start
+4. Entropy-Based Early Stopping:
+  • AGR halts expansion when confidence exceeds threshold
+  • Prevents wasteful over-exploration
+  • Average nodes retrieved: 127 (vs 500+ for flat top-k)
+  • Token reduction: 65% compared to context-stuffing approaches
+5. Vertical Integration Benefits:
+  • Full control over the debugging pipeline without external dependencies
+  • Optimized end-to-end: retriever $\rightarrow$ LLM $\rightarrow$ validator
+  • Shared embeddings between PDM and AGR reduce redundant computation
+  • Custom optimization for debugging-specific workloads
+AGR Efficiency Summary Key Efficiency Gains: Graph traversal optimized for debugging patterns Significantly reduced token usage (31.2K avg vs 89K+ for competitors) Containerized test execution for rapid validation Incremental memory updates avoid redundant processing Result: More efficient debugging at scale compared to API-based approaches
+TABLE V: Why Chronos Can Run AGR Cheaply: Architectural Advantages
+Advantage How Chronos Achieves It Why Others Can’t
+Narrow scope Focused on bug repair only Others support all codegen tasks
+3.7 avg retrievals per debug 15+ retrievals for autocomplete
+Smart retrieval Graph traversal + entropy-based stopping Others use flat top-k retrieval
+Halts at 89% confidence threshold Fixed k regardless of confidence
+Shared graph Built once per repo, updated incrementally Others re-embed on every query
+Extremely low amortized cost High per-query embedding cost
+Stateful memory PDM learns across debugging sessions Others are stateless, start fresh
+87% cache hit on recurring patterns 0% reuse, full retrieval each time
+Vertical stack Own retriever + LLM + test runner Others rely on external APIs
+Integrated efficiency gains API costs with lower success rates
+III-F Multi-Code Compositional Retrieval: Beyond Token Windows
+Chronos implements a sophisticated retrieval mechanism that goes beyond simple embedding similarity to understand the complex relationships in debugging contexts.
+III-F1 Multi-Code Compositional Retrieval
+Upon each debugging request, the Adaptive Retrieval Engine builds a focused context window through:
   • Issuing semantic queries to the Memory Engine that leverage both metric similarity and structural navigation in the code graph, with dynamic depth expansion (k-hop) based on query complexity.
@@
   • Utilizing edge type priorities: implementation edges (weight=1.0), dependency edges (weight=0.8), documentation edges (weight=0.6), ensuring most relevant paths are explored first.
-TABLE II: Example multi-code association retrieval: constructing a task-specific context window for a bug fix.
+TABLE VI: Example multi-code association retrieval: constructing a task-specific context window for a bug fix.
 Step Retrieved Entity Relationship
@@
 Q5 commit_a1b2c3 Last related commit
-This approach allows Chronos to reason across arbitrarily distant, compositionally linked code and documentation artifacts—precisely what is needed for complex debugging, cross-module dependencies, or audit trails.
-III-F Adaptive Graph-Guided Retrieval (AGR)
+This approach allows Chronos to reason across arbitrarily distant, compositionally linked code and documentation artifacts, precisely what is needed for complex debugging, cross-module dependencies, or audit trails.
+III-G Adaptive Graph-Guided Retrieval (AGR)
 Traditional flat retrieval approaches fail to capture the intricate relationships between code artifacts, leading to incomplete context and erroneous fixes.
 Chronos introduces Adaptive Graph-Guided Retrieval (AGR), a dynamic mechanism that intelligently expands retrieval neighborhoods based on query complexity and confidence thresholds.
-III-F1 Iterative Context Expansion
+III-G1 Comparison with State-of-the-Art RAG Techniques
+The 2025 landscape of RAG techniques has evolved significantly, yet each approach faces limitations when applied to debugging:
+HyDE (Hypothetical Document Embeddings) [15] generates synthetic answers to improve retrieval but struggles with debugging where incorrect hypotheses can mislead the search process.
+While HyDE achieves 42% improvement in general retrieval tasks (measured on MS MARCO), it shows only 8% improvement for debugging scenarios on our MRR benchmark.
+Self-RAG [25] uses reflection tokens (ISREL, ISUSE) to dynamically decide retrieval necessity.
+However, debugging requires continuous retrieval across multiple hops, making binary retrieval decisions insufficient.
+Self-RAG achieves 31% debug success rate on MRR benchmark when combined with GPT-4.1.
+FLARE (Forward-Looking Active Retrieval) [16] monitors generation confidence to trigger retrieval.
+This works well for sequential text generation but fails in debugging where confidence doesn’t correlate with correctness, models can be confidently wrong about bug fixes.
+Graph RAG [26] integrates knowledge graphs but typically uses static graphs that don’t capture the dynamic nature of evolving codebases.
+Standard Graph RAG achieves 28% success rate on cross-file debugging tasks in the MRR benchmark.
+In contrast, Chronos’s AGR combines the benefits of these approaches while addressing their limitations through adaptive k-hop expansion, typed edge traversal, and confidence-based termination specifically tuned for debugging workflows.
+III-G2 Why Graph Traversal Outperforms Naive Chunked Retrieval
+Consider a concrete example that illustrates the fundamental limitation of linear retrieval.
+A null pointer exception occurs in PaymentProcessor.java:142 when processing refunds.
+The stack trace shows:
+java.lang.NullPointerException
+at PaymentProcessor.processRefund(PaymentProcessor.java:142)
+at RefundService.handleReturn(RefundService.java:89)
+at OrderController.cancelOrder(OrderController.java:234)
+Naive chunked retrieval would:
+  1. 1.
+    Retrieve top-k chunks around line 142 based on embedding similarity
+  2. 2.
+    Miss that customerAccount is null because initialization happens in AccountService.java
+  3. 3.
+    Fail to connect that recent commit changed config/payment.yml timeout from 30s to 5s
+  4. 4.
+    Generate a band-aid null check instead of fixing the root cause
+AGR’s graph traversal instead:
+  1. 1.
+    Starts at error location (PaymentProcessor:142)
+  2. 2.
+    Follows data flow edge: customerAccount $\leftarrow$ AccountService.loadAccount()
+  3. 3.
+    Follows temporal edge: payment.yml modified 2 days ago
+  4. 4.
+    Discovers timeout now expires before account loads, causing null
+  5. 5.
+    Proposes correct fix: adjust timeout or add async handling
+This demonstrates AGR’s key advantage: it follows causal paths rather than textual similarity, achieving O(k·d) complexity where k=hops and d=average degree, versus O(n) for naive retrieval over n chunks.
+III-G3 Building on Existing Graph-Based Code Understanding
+While AGR represents a unified breakthrough, it builds upon fragmented components that exist across various systems. Understanding this landscape helps position AGR’s contributions:
+1. Graph-Style Retrieval in Current Systems
+Several modern systems employ basic graph structures for code understanding.
+Top SWE-bench submissions like Magicoder-S with retrieval agents construct simple import trees and dependency graphs.
+SWE-agent pioneered using test files to backtrack to relevant functions, establishing the value of test-to-code linkage.
+However, these implementations remain primitive:
+  • Graphs are shallow, typically exploring only 1-2 hops from the starting point
+  • No confidence-aware traversal - they retrieve everything within a fixed radius
+  • Missing temporal dimensions - no commit history or evolution tracking
+  • Lack causal chains connecting logs $\rightarrow$ stack traces $\rightarrow$ code $\rightarrow$ tests $\rightarrow$ PRs
+  • No adaptive halting based on information sufficiency
+What these systems demonstrate is that graph construction and basic traversals are valuable for code understanding.
+AGR takes this foundation and adds the depth control, edge weighting, and test-memory integration needed for effective debugging.
+2. Multi-Hop Retrieval from NLP Research
+The NLP community has explored multi-hop reasoning extensively.
+Models like DR-BERT [45] perform multi-hop retrieval for question answering by following entity links across Wikipedia.
+REALM [46] showed that retrieval-augmented pretraining improves downstream tasks.
+GNN-RAG [47] combines graph neural networks with retrieval for knowledge-intensive tasks.
+These approaches inspired code-specific adaptations like RAG-SweBench+, which attempts multi-hop retrieval in codebases. However, fundamental differences limit their debugging effectiveness:
+  • Text-based retrieval misses code structure (AST, type systems, execution flow)
+  • No integration with debugging artifacts (logs, stack traces, test failures)
+  • Designed for factual QA, not causal reasoning about bug origins
+  • Lack domain-specific edge types (imports, inherits, calls, emits-log)
+AGR adapts the multi-hop principle but grounds it in code-specific structures and debugging workflows.
+3. Memory Systems in Code Agents
+Projects like Magicoder-S-Agent, Octocoder, and CodeAgent maintain various forms of persistent memory.
+Some store chunk embeddings across sessions, while others cache test outcomes and past fix attempts.
+These systems prove that memory helps, but their retrieval remains simplistic:
+  • Flat top-k retrieval from memory banks, no graph traversal
+  • No semantic paths from failure signals through memory
+  • Memory updates are append-only, not integrated with confidence scores
+  • Cannot trace causal chains through historical fixes
+AGR’s integration with PDM enables true graph-guided memory retrieval, where past debugging sessions inform current traversal paths.
+4. Internal Tools at Tech Companies
+Major tech companies have built sophisticated internal systems that parse monorepos into searchable graphs.
+Google’s internal code search constructs AST and dependency graphs across billions of lines.
+Meta’s code understanding tools link stack traces to potential causes.
+Sourcegraph provides semantic code navigation across repositories.
+While these tools are powerful, they remain:
+  • Proprietary and inaccessible to researchers
+  • Focused on human-assisted search, not autonomous debugging
+  • Lacking integration with language models for fix generation
+  • Missing the iterative test-fix-refine loops needed for debugging
+AGR’s Unified Contribution
+What makes AGR unique is not any single component, but the unified system that combines:
+  1. 1.
+    Multi-signal graph construction: AST + logs + tests + PRs + commits in one traversable structure
+  2. 2.
+    Confidence-aware adaptive expansion: Exploring exactly as deep as needed, no more, no less
+  3. 3.
+    Memory-integrated traversal: Learning from past debugging sessions to guide future paths
+  4. 4.
+    Debugging-specific optimization: Trained on 15M real debugging scenarios, not general retrieval
+  5. 5.
+    Autonomous operation: Full integration with test loops and fix validation
+By unifying these fragmented approaches and adding debugging-specific innovations, AGR achieves the 4-5x performance improvement over existing systems.
+While individual techniques used in AGR, such as multi-hop retrieval, AST graph parsing, and chunk memory, have appeared in prior work, Chronos is the first system to unify these methods into a debugging-first retrieval engine.
+AGR combines graph-guided, edge-weighted traversal with confidence-aware stopping, memory integration, and dynamic context composition explicitly tuned for root-cause tracing and patch generation.
+TABLE VII: AGR Components: Fragmented Existence vs Unified in Chronos
+Capability Exists Elsewhere? Chronos
+AST + Log + PR + Test graph Fragmented ✓ Unified graph
+Confidence-aware hop limit No
+Graph-guided memory integration No
+Retrieval tuned for debugging No
+Multi-turn patch validation Rare ✓ Core loop
+This positions Chronos as the first unified, publicly described system that brings together all the pieces needed for effective autonomous debugging at repository scale.
+3) Limitations of Orchestration Frameworks in Debugging:
+Popular orchestration frameworks like LangChain and LangGraph, while powerful for general AI applications, face fundamental limitations when applied to debugging:
+LangChain’s Context Fragmentation: LangChain’s chain-based approach treats each step as independent, leading to context loss between debugging iterations.
+When combined with GPT-4.1, it achieves only 18% debug success due to: - Stateless chains that cannot maintain debugging history across iterations - Message history overflow when debugging sessions exceed context limits - Chain-of-thought breaking when bugs require backtracking or re-evaluation
+LangGraph’s Graph Limitations: Despite having graph-based memory, LangGraph achieves only 22% debug success because: - Static graph traversal cannot adapt to the iterative nature of debugging - Node-based processing loses fine-grained code relationships - State persistence fails when debugging requires cross-session memory
+Chain-of-Thought Degradation: Traditional prompting techniques degrade severely on debugging tasks: - CoT prompting with Claude 4 Opus: drops from 92.8% (code generation) to 15% (debugging) - ReAct with GPT-4.1: 17% debug success despite structured reasoning - Tree-of-Thoughts: 19% as exploration trees explode with multi-file dependencies
+These frameworks excel at sequential tasks but fundamentally misunderstand debugging’s need for persistent memory, iterative refinement, and cross-session learning, capabilities that Chronos provides natively.
+III-G4 Iterative Context Expansion
 The AGR mechanism operates through iterative k-hop neighbor expansion:
@@
     Confidence-Based Termination: Stop expansion when retrieval confidence exceeds threshold or diminishing returns detected
-III-F2 Graph-Guided vs Traditional Planning
+5) AGR Algorithm Details:
+The following algorithm formally describes the Adaptive Graph-Guided Retrieval process:
+Algorithm 1 Adaptive Graph-Guided Retrieval (AGR)
+1:Query $q$, Code Graph $G=(V,E)$, Confidence threshold $\tau$
+2:Retrieved context $C$
+3$\triangleright$ Initialize
+[…]
+[…]
+[…]
+[…]
+8$\triangleright$ Adaptive expansion
+9:while $\textsc{Confidence}(C,q)<\tau$ and $k\leq k_{max}$ do
+10:  $candidates\leftarrow\emptyset$
+11:  for $node\in seeds$ do
+[…]
+13:   for $n\in neighbors\setminus visited$ do
+[…]
+15:     $candidates\leftarrow candidates\cup\{(n,score)\}$
+16:   end for
+17:  end for
+18$\triangleright$ Select top candidates based on typed edges
+[…]
+20:  for $(node,score)\in selected$ do
+[…]
+[…]
+23:     $visited\leftarrow visited\cup\{node\}$
+24:   end if
+25:  end for
+26$\triangleright$ Adaptive depth adjustment
+27:  if $\textsc{DeltaConfidence}(C)<\epsilon$ then
+28:   $k\leftarrow k+1$triangleright Expand search radius
+29:  end if
+[…]
+31:end while
+32:
+33:return $C$
+Key innovations in the AGR algorithm (Algorithm 1):
+  • Dynamic k-hop adjustment: Unlike static retrieval depths, AGR adaptively increases $k$ based on confidence improvements
+  • Typed edge prioritization: Implementation and dependency edges receive higher weights than generic references
+  • Semantic seed extraction: Initial nodes are selected based on deep semantic understanding, not just keyword matching
+  • Confidence-driven termination: Retrieval stops when sufficient context is gathered, avoiding noise from over-retrieval
+6) Theoretical Analysis of AGR:
+[…]
+Theorem 1 (AGR Retrieval Complexity)
+[…]
+Proof:
+At each iteration $k$, we explore at most $|S|\cdot d^{k}$ nodes.
+Sorting candidates requires $O(d^{k}\log(d^{k}))$ time.
+The algorithm terminates when confidence exceeds $\tau$ or $k=k_{max}$, giving the stated bound.
+∎
+Convergence Properties:
+Theorem 2 (Confidence Convergence)
+Under the assumption that relevance scores follow a power-law distribution with exponent $\alpha>1$, the confidence function $\mathcal{C}(C,q)$ converges to a value $c^{*}\geq\tau$ with probability $1-\delta$ after $k^{*}=O(\log_{d}(1/\delta))$ iterations.
+Proof:
+The confidence function is defined as:
+$\mathcal{C}(C,q)=1-H(C|q)/H_{max}$
+where $H(C|q)$ is the conditional entropy.
+As we add relevant nodes, entropy decreases monotonically.
+Under power-law distribution, most relevant nodes are within $O(\log_{d}n)$ hops, ensuring convergence.
+∎
+Bounded Retrieval Path Cost:
+Lemma 1 (Path Cost Bound)
+The total retrieval cost is bounded by $O(|S|\cdot\lambda\cdot k_{max}^{2}\cdot d^{k_{max}})$ where $\lambda$ is the selection ratio per hop.
+This theoretical foundation ensures AGR’s efficiency even on large codebases with millions of nodes.
+Queryauth.pytest.pyuser.pydb.pyconfigapi.pycachelogsdocsissueutil
+Confidence
+k=1: 45%
+k=2: 78%
+k=3: 92% $>\tau$
+Seeds (k=1) Retrieved (k=2)
+Candidates (k=3) Not explored
+StartStop: conf $>\tau$
+Figure 6: Adaptive Graph-Guided Retrieval (AGR) visualization.
+The algorithm starts from a query, extracts semantic seed nodes, and iteratively expands the retrieval neighborhood (k-hops) until confidence exceeds threshold $\tau$.
+Edge types and relevance scores guide the expansion process.
+Key Takeaways: AGR Algorithm Dynamic Expansion: k-hop depth adapts based on query complexity and confidence Typed Traversal: Implementation and dependency edges prioritized over generic links Early Termination: Stops when confidence exceeds threshold, avoiding over-retrieval Performance: Achieves 92% precision at 85% recall on debugging queries
+III-G5 Graph-Guided vs Traditional Planning
 Our empirical analysis reveals fundamental differences between traditional LLM planning and AGR-enhanced debugging:
-[…]
+Traditional LLM PlanningQuery: Implement state machinestate <= `d<0> > SS () <-d<0> > S Traditional Steps:1.
+Define Module Interface2.
+Define State Encoding3.
+State Transition Logic4.
+Output Logic: Assign outputsIssues:- High-level plans without task details- Hard to follow implementation- Lost signal/transition specs Generated Code:assign S_next = (state == 5’b11010) ?5’b11010 : (state == 5’b10110) ?5’b11010 : state;$\times$ Incorrect implementation AGR-Enhanced DebuggingQuery: Implement state machine+ Graph retrieval of specifications+ Signal transition examples AGR-Guided Steps:1.
+Retrieve signal definitions (k=1)2.
+Expand to transitions (k=2)3.
+Include test examples (k=3)Retrieved Context:- S1_next: Output signal- Wait->S, S->S transitions- Example: 9’b101000100 Generated Code:// Correct implementationassign S1_next = S;// Based on retrieved $\checkmark$ Verified correct 23% Success Rate87% Success Rate
 Figure 7: Traditional LLM planning vs AGR-enhanced debugging: Graph-guided retrieval provides complete context, leading to accurate implementations.
@@
 Figure 9: Multi-modal retrieval mechanism in Chronos.
-III-G Reasoning, Generation, and Autonomous Orchestration
+III-H From Reasoning to Resolution: Autonomous Debug Orchestration
 The transformer-based Chronos Reasoning Model operates directly over the retrieved, multi-source debugging context. Unlike classical code completion models, Chronos:
@@
   • Produces stepwise fix plans, code diffs, documentation updates, and regression test suggestions in a unified, automated debugging loop.
-  • Orchestrates a full debugging workflow: proposes bug fixes, invokes relevant tests, parses results, iterates on failures, and generates changelogs or PR summaries—all autonomously.
+  • Orchestrates a full debugging workflow: proposes bug fixes, invokes relevant tests, parses results, iterates on failures, and generates changelogs or PR summaries, all autonomously.
 All outputs and feedback streams (test results, reviewer comments, CI/CD events) are fed back into the Memory Engine, enabling lifelong refinement and rapid adaptation to new debugging scenarios.
@@
 Figure 10: Chronos debugging feedback loop: Automated bug fix generation, validation, plan refinement, and memory update for continuous autonomous improvement.
-IV THE DEBUGGING LANGUAGE MODEL PARADIGM
-Kodezi Chronos fundamentally departs from traditional code models by being purpose-built as a debugging language model—the first of its kind.
+IV DEBUGGING AS A DISTINCT ML TASK: THE CHRONOS PARADIGM
+Kodezi Chronos fundamentally departs from traditional code models by being purpose-built as a debugging language model, the first of its kind.
 While existing LLMs treat debugging as a code generation problem, Chronos recognizes it as a complex, iterative process requiring specialized capabilities, training, and architecture.
-IV-A Core Architectural Layers
-Chronos implements a revolutionary 7-layer architecture specifically designed for autonomous debugging:
+IV-A Seven-Layer Debugging Architecture: Specialized Components
+Chronos implements a 7-layer architecture designed for autonomous debugging (illustrated in Figure 11):
   1. 1.
-    Multi-Source Input Layer: Ingests heterogeneous debugging signals including source code, CI/CD logs, error traces, stack dumps, configuration files, historical PRs, and issue reports.
+    Multi-Source Input Layer: Ingests diverse debugging inputs including source code, CI/CD logs, error traces, stack dumps, configuration files, historical PRs, and issue reports.
 Unlike code models that primarily process source files, Chronos natively understands debugging artifacts.
@@
     • Risk assessment reports
-IV-B Training on Debugging Workflows
+1. Multi-Source Input2. Adaptive Retrieval (AGR)3. Debug-Tuned LLM4. Orchestration5. Persistent Memory6. Execution Sandbox7. ExplainabilityIterateUpdate
+Figure 11: The 7-layer architecture of Chronos. Each layer is specialized for debugging tasks, with bidirectional information flow enabling iterative refinement and continuous learning.
+IV-B Debug-Specific Training: 15M Real-World Bug Scenarios
 Unlike models trained primarily on code completion, Chronos’s training regime focuses exclusively on debugging scenarios:
@@
   • Cross-Repository Pattern Recognition: Identifying similar bugs across different codebases
-IV-C The Autonomous Debugging Loop
-Chronos’s debugging loop represents a fundamental innovation over single-shot code generation:
+IV-C Autonomous Fix-Test-Refine Loop: Iterative Convergence
+Chronos’s debugging loop represents a fundamental innovation over single-shot code generation. Algorithm 2 presents the core logic:
+Algorithm 2 Fix-Test-Refine Loop
+1:Bug report $B$, Codebase $C$, Test suite $T$, PDM memory $M$
+2:Validated fix $F^{*}$ or failure report
+3:
+[…]
+[…]
+[…]
+7:
+8:while $k<$ MAX_ITERATIONS do
+9:  $F_{k}\leftarrow\text{Chronos.propose\_fix}(B,context,patterns)$
+10:  $result\leftarrow\text{Sandbox.execute}(F_{k},T)$
+11:  if $result$.success then
+12:   $regression\leftarrow\text{Sandbox.run\_extended\_tests}(F_{k})$
+13:   if $\neg regression$ then
+[…]
+15:     return $F_{k}$ as $F^{*}$
+16:   end if
+17:  end if
+18:  $context\leftarrow context\cup\text{Analyzer.extract\_failure}(result)$
+19:  $patterns\leftarrow patterns\cup\text{PDM.similar\_failures}(result)$
+20:  $k\leftarrow k+1$
+21:end while
+22:
+23:return ”Failed to converge after {$k$} iterations”
+The key innovation is the feedback loop: each failed attempt enriches the context with failure analysis, making subsequent attempts more informed:
 Detect IssueRetrieve ContextPropose FixRun TestsTests Pass?Commit & DeployRefine StrategyUpdate MemoryYesNo
@@
 This loop continues autonomously, with each iteration informed by previous attempts and accumulated knowledge, until a validated fix is achieved or human intervention is requested.
-V EVALUATION METHODOLOGY
+Key Takeaways: Autonomous Debugging Loop Iterative Refinement: Unlike single-shot generation, continuously improves fixes based on test results Memory Integration: Each iteration learns from previous attempts, avoiding repeated failures Autonomous Operation: Requires no human intervention for 65.3% of real-world bugs Efficiency: Average 2.2 iterations to successful fix vs 4.8 for competing systems
+IV-D Runtime Execution Analysis: Latency and Flow Dynamics
+To better understand Chronos’s runtime behavior, we present a detailed flow diagram showing the actual execution path during a debugging session:
+Bug Report/Error SignalAGR Retrieval(k-hop expansion)Code Context+ HistoryGenerate FixHypothesisApply CodeChangesRun TestsCommit FixUpdate MemoryAnalyzeFailureBacktrack?RefineHypothesis12ms178ms523ms89ms1.2sPassFail156msYesNoAvg: 2.2 iterationsto successMemory updateon each iteration
+Figure 13: Runtime execution flow showing typical latencies. The fix-loop iterates autonomously until tests pass or confidence threshold is exceeded.
+V EVALUATION: COMPREHENSIVE DEBUGGING BENCHMARKS
 To rigorously assess Kodezi Chronos’s capabilities across realistic debugging and maintenance workflows, we adopt a multi-faceted evaluation strategy that goes beyond conventional sequence completion or shallow retrieval tests.
-V-A Benchmarking Datasets and Tasks
-Chronos is evaluated on a diverse suite of benchmarks, comprising:
-  • Standard Code Generation Benchmarks: HumanEval [2], MBPP [3], and related tasks for basic code synthesis and function-level reasoning.
-  • Repository-Scale Debugging Tasks: A curated set of real-world bug reports, failing test scenarios, and codebase refactoring challenges drawn from open-source projects and industry case studies.
-  • Multi-Code Association Retrieval: Custom synthetic benchmarks requiring retrieval and reasoning across arbitrarily placed, subtly linked code/documentation artifacts—simulating variable tracing, cross-file bug localization, or dependency update chains.
-V-B Multi-Code Reasoning Evaluation Protocol
+05005005001,0001{,}0001 , 0001,5001{,}5001 , 5002,0002{,}0002 , 0002,5002{,}5002 , 5003,0003{,}0003 , 0003,5003{,}5003 , 5004,0004{,}0004 , 0004,5004{,}5004 , 5005,0005{,}0005 , 0005,5005{,}5005 , 500 Simple Bug (1-2 files) Medium Bug (3-5 files) Complex Bug (6-10 files) Cross-Module (10+ files) Historical Bug (w/ commits) 5.18s4.22s2.68s1.68s900msRetrieval Time (ms)Graph ConstructionSemantic Searchk-hop TraversalContext Assembly
+Figure 14: AGR retrieval time breakdown by bug complexity.
+Even for complex cross-module bugs requiring 10+ file analysis, total retrieval completes in under 5.2 seconds, enabling rapid debugging iterations.
+V-A Statistical Methodology
+All experiments follow rigorous statistical protocols to ensure reproducibility and validity:
+Power Analysis: Sample sizes determined via G*Power 3.1 to achieve 0.95 power at $\alpha=0.05$ for detecting Cohen’s $d\geq 0.8$ effect sizes.
+Minimum N=64 per condition for between-subjects comparisons.
+Effect Size Measures:
+  • Cohen’s d for continuous outcomes: $d=\frac{\mu_{1}-\mu_{2}}{\sigma_{pooled}}$
+  • Cliff’s delta for ordinal data: $\delta=\frac{2U}{n_{1}n_{2}}-1$
+  • Cramér’s V for categorical comparisons: $V=\sqrt{\frac{\chi^{2}}{n(k-1)}}$
+Statistical Tests:
+  • Wilcoxon signed-rank test for paired comparisons (non-parametric)
+  • Mann-Whitney U test for independent samples
+  • Bonferroni correction for multiple comparisons: $\alpha_{adj}=\frac{0.05}{m}$
+  • Bootstrap confidence intervals (10,000 resamples) for all metrics
+Inter-rater Reliability: Bug classification achieved Fleiss’ $\kappa=0.87$ (substantial agreement) across 3 expert annotators on 500 randomly sampled bugs.
+[…]
+Cross-Validation Protocol:
+  • 5-fold stratified cross-validation: Stratified by bug type, language, and complexity
+  • Repository-level splits: Entire repositories assigned to either train or test to prevent data leakage
+  • Temporal validation: Additional holdout set with bugs from 2025 (post-training)
+  • Nested CV for hyperparameters: Inner 3-fold CV for AGR threshold tuning ($\tau\in[0.7,0.95]$)
+Variance Reporting: All metrics reported as mean ± standard deviation across CV folds, with 95% confidence intervals.
+V-B Evaluation Framework and Datasets
+V-B1 Dataset Sources and Composition
+Chronos is evaluated on a comprehensive suite of benchmarks:
+  1. 1.
+    Public Debugging Datasets:
+    • Defects4J: 438 real bugs from 17 Java projects including Apache Commons, JFreeChart, and Closure Compiler
+    • BugsInPy: 493 bugs from 17 Python projects including pandas, keras, and tornado
+    • SWE-bench++: 2,294 GitHub issues requiring repository-wide changes, extended with test harnesses
+    • COAST: 283 multi-file debugging scenarios from open-source C/C++ projects
+    • MLDebugging: 412 machine learning bugs including tensor shape mismatches and gradient issues
+    • MdEval: 196 memory corruption and data race bugs from systems software
+  2. 2.
+    Proprietary Enterprise Dataset (Anonymized):
+    • 3,842 debugging sessions from Fortune 500 companies (with signed data agreements)
+    • All code snippets anonymized using differential privacy techniques ($\epsilon=1.2$)
+    • Sensitive identifiers replaced with semantic placeholders preserving debugging context
+    • Manual review by security team to ensure no IP leakage
+  3. 3.
+    Synthetic Bug Generation:
+    • 8,000 synthetically injected bugs using mutation testing frameworks
+    • Bug types: null pointer exceptions, off-by-one errors, race conditions, API misuse
+    • Generated from top 1,000 GitHub repositories across 12 languages
+    • Each bug verified to compile and fail exactly one test
+V-C Baseline Systems
+We compare Chronos against state-of-the-art models and specialized debugging tools:
+General-Purpose LLMs:
+  • GPT-4.1 (OpenAI): 1.8T parameters, 128K context, with custom debugging prompt
+  • Claude 4 Opus (Anthropic): 200K context, strongest on code understanding
+  • Gemini 2.5 Pro (Google): 2M context window, multimodal capabilities
+  • DeepSeek V3 (DeepSeek): 671B MoE, 37B active, cost-efficient
+Specialized Debugging Tools:
+  • Microsoft IntelliCode Compose: IDE-integrated debugging suggestions
+  • DeepDebug (Microsoft Research): Neural bug localization system
+  • SWE-agent + GPT-4: Agentic debugging with 22.9% SWE-bench
+  • AutoCodeRover: Structure-aware debugging, 19.7% on SWE-bench
+  • LangGraph + ReAct + Claude 4: Multi-agent debugging pipeline
+Enhanced RAG Baselines:
+  • GPT-4.1 + HyDE: Hypothetical document embeddings for retrieval
+  • Claude 4 + Self-RAG: Self-reflective retrieval augmentation
+  • Gemini + GraphRAG: Knowledge graph enhanced retrieval
+CI/CD Integration Data:
+  • 15M+ CI/CD logs from public GitHub Actions and Jenkins builds
+  • Stack traces and error logs extracted with automated parsers
+  • Build failure patterns categorized into 127 common root causes
+  • Privacy preserved by filtering any URLs, credentials, or personal data
+V-D Debugging-Specific Benchmarks
+Beyond general code generation benchmarks like HumanEval and MBPP [51], we evaluate Chronos on specialized debugging benchmarks that better reflect real-world maintenance challenges:
+V-D1 COAST (Code Optimization and Analysis for Software Teams)
+COAST [52] evaluates debugging through 5,000 production bug scenarios requiring:
+  • Cross-repository dependency tracking
+  • Performance regression identification
+  • Security vulnerability patching
+  • API migration debugging
+V-D2 MLDebugging Benchmark
+MLDebugging [53] focuses on machine learning pipeline failures with 2,500 cases covering:
+  • Training instability diagnosis
+  • Data pipeline corruption detection
+  • Model serving failures
+  • Gradient explosion/vanishing debugging
+V-D3 MdEval (Multi-dimensional Evaluation)
+MdEval [54] provides comprehensive debugging evaluation across:
+  • Concurrency bugs (race conditions, deadlocks)
+  • Memory management issues (leaks, corruption)
+  • Network protocol violations
+  • Build system configuration errors
+TABLE VIII: Consolidated debugging benchmark performance.
+Benchmark Task Category Chronos Amazon Q ACR Claude 4 Opus GPT-4.1
+COAST Cross-repository deps 71.2 43.8 31.4 18.2 15.7
+Performance bugs 68.4 41.2 28.7 14.3 12.8
+Security patches 74.8 48.3 35.2 21.1 19.4
+API migration 69.7 45.1 32.8 17.6 16.2
+Average 71.0 44.6 32.0 17.8 16.0
+MLDebug Training instability 65.3 38.7 27.1 11.8 10.2
+Data pipeline 72.1 44.2 31.8 15.4 13.7
+Model serving 67.8 41.3 29.4 13.2 11.9
+Gradient issues 61.4 35.8 24.3 9.7 8.4
+Average 66.7 40.0 28.2 12.5 11.1
+MdEval Concurrency 58.7 32.1 22.8 7.3 6.1
+Memory mgmt 63.2 36.4 25.7 10.2 8.8
+Network bugs 66.9 40.8 28.3 12.7 11.4
+Build config 70.5 42.7 30.1 16.8 14.9
+Average 64.8 38.0 26.7 11.8 10.3
+Overall Average 67.5*** 41.7 29.0 14.1 12.4
+ACR = AutoCodeRover. ***p <0.001<0.001< 0.001 compared to best baseline (Amazon Q), two-tailed t-test, n=12,500
+V-E Multi-Code Reasoning Evaluation Protocol
 Unlike traditional benchmarks that target token-level prediction in narrow context, our protocol explicitly:
@@
   • Measures both retrieval accuracy (whether Chronos finds all necessary context) and end-to-end task success (whether it can autonomously fix, validate, and document the issue).
-V-C Multi Random Retrieval Benchmark
-We introduce the Multi Random Retrieval (MRR) benchmark, specifically designed to evaluate debugging-oriented retrieval capabilities:
-V-C1 Benchmark Design
+V-F Multi Random Retrieval Benchmark
+We introduce the Multi Random Retrieval (MRR) benchmark, specifically designed to evaluate debugging-oriented retrieval capabilities. The full evaluation suite is scheduled for release in Q1 2026.
+V-F1 MRR Design and Methodology
+The Multi Random Retrieval benchmark addresses a critical gap in existing evaluations: real-world debugging requires finding scattered information across large codebases where traditional similarity-based retrieval fails.
+Key Design Principles:
+  1. 1.
+    Random Distribution: Relevant debugging clues are randomly scattered across 10-50 files, simulating real-world information dispersion
+  2. 2.
+    High Noise Ratio: 70% of retrievable content is plausible but irrelevant, testing precision
+  3. 3.
+    Multi-Hop Requirements: Average 3-7 retrieval steps needed to gather complete context
+  4. 4.
+    Temporal Scattering: Information spans multiple commits/time periods
+Formal Metrics:
+  • Precision@k: Fraction of retrieved chunks that are actually relevant to the bug fix
+  • Recall@k: Fraction of all relevant chunks successfully retrieved
+  • MRR Score: Mean reciprocal rank of first correct retrieval, computed as $\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{rank}_{i}}$
+  • Fix Success Rate: Whether the generated fix actually resolves the bug
+Bug: Async Task FailureTaskRunner.javaAsyncUtils.javaErrorHandler.javaWrong Fix: Add try-catchText similarityTaskExecutor.java:89ThreadPoolConfig.ymlcommit: ”reduce threads”Correct Fix: Restore pool sizeStack traceConfig refGit blameGPT-4: Surface Pattern MatchingChronos: Causal Path Following
+Figure 15: MRR example showing retrieval paths: GPT-4 follows textual similarity to related but incorrect files, while Chronos traces causal dependencies through stack trace → configuration → git history to find the true root cause.
+V-F2 Reproducibility and Open Science
+To ensure reproducibility and enable fair comparisons, we provide:
+  1. 1.
+    Open Evaluation Subset:
+    • 500 debugging scenarios (10% of full benchmark) with complete test harnesses
+    • Ground truth fixes and intermediate retrieval annotations
+    • Automated evaluation scripts with standardized metrics
+    • Docker containers with exact environment configurations
+  2. 2.
+    Evaluation Infrastructure:
+    • chronos-eval: Python package for running benchmarks
+    • Supports custom model integration via simple API
+    • Automated result validation and statistical significance testing
+    • Leaderboard submission system with blind test set
+  3. 3.
+    Detailed Task Definitions:
+Example MRR Task Format:
+"bug_id": "apache-commons-math-1234",
+"repo_snapshot": "git://github.com/apache/commons-math@abc123",
+"failing_tests": ["TestLinearRegression.testFit"],
+"scattered_files": [
+"src/.../stat/regression/AbstractRegression.java",
+"src/.../linear/MatrixUtils.java"
+// ... 15 more files across different modules
+"temporal_range": "2023-01-15 to 2023-04-22",
+"evaluation": {
+"fix_validator": "docker run chronos-eval:fix-validator",
+"context_scorer": "weighted_jaccard",
+"regression_tests": ["TestSuite.class"]
+V-F3 Benchmark Design
 The MRR benchmark consists of 5,000 real-world debugging scenarios where:
@@
     Multi-Modal Artifacts: Solutions require combining code, tests, logs, and documentation
-V-C2 Evaluation Metrics
+V-F4 Evaluation Metrics
   • Retrieval Precision@k: Fraction of retrieved artifacts that are relevant to the bug fix
@@
   • Context Efficiency: Ratio of used vs retrieved tokens in the final solution
-V-C3 Results on MRR Benchmark
-TABLE III: Performance on Multi Random Retrieval benchmark, demonstrating Chronos’s superior ability to find and utilize scattered debugging context.
+V-F5 Results on MRR Benchmark
+TABLE IX: Performance on Multi Random Retrieval benchmark, demonstrating Chronos’s superior ability to find and utilize scattered debugging context.
 Model Precision@10 Recall@10 Fix Accuracy Context Eff.
-GPT-4 + RAG 42.3% 31.7% 8.9% 0.23
-Claude-3 + Vector DB 48.1% 36.2% 11.2% 0.28
-Gemini-1.5 + Graph 51.7% 41.8% 14.6% 0.31
+GPT-4.1 + RAG 55.2% 42.3% 13.8% 0.34
+Claude 4 Opus + RAG 62.1% 48.7% 14.2% 0.41
+Gemini 2.5 Pro + RAG 51.7% 40.1% 12.4% 0.38
 Kodezi Chronos 89.2% 84.7% 67.3% 0.71
-V-D Adaptive Graph-Guided Retrieval Performance
+V-F6 Comprehensive MRR Results Against State-of-the-Art Debugging Systems
+The MRR benchmark reveals the fundamental difference between general-purpose code models and debugging-specific systems. Table X shows detailed performance across all evaluated systems:
+TABLE X: Comprehensive MRR benchmark results comparing Chronos against modern debugging systems and general-purpose code models.
+Results demonstrate the critical importance of debugging-specific design.
+System Type Precision@10 Recall@10 Fix Acc. Iterations Success Time
+General-Purpose Code Models (2025)
+Claude 4 Opus Code Gen 62.1% 48.7% 14.2% 2.3 15.2 min
+Claude 4 Sonnet Code Gen 59.8% 46.3% 13.1% 2.1 14.8 min
+GPT-4.1 Code Gen 55.2% 42.3% 13.8% 1.8 12.3 min
+GPT-4o Code Gen 48.3% 37.2% 10.2% 1.6 11.7 min
+Gemini 2.5 Pro Code Gen 51.7% 40.1% 12.4% 2.0 13.5 min
+DeepSeek V3 Code Gen 44.2% 34.8% 9.7% 1.4 10.2 min
+Qwen2.5-Coder-32B Code Gen 41.3% 31.2% 8.3% 1.3 9.8 min
+IDE-Integrated Systems
+Cursor Agent Mode IDE 32.4% 24.1% 4.2% 1.0 8.5 min
+Windsurf Cascade IDE 35.7% 27.3% 5.1% 1.2 9.1 min
+Claude Code CLI CLI 40.2% 31.8% 6.8% 1.4 10.3 min
+Gemini CLI (1.5 Pro) CLI 43.1% 35.2% 9.7% 1.5 11.2 min
+Debugging-Focused Systems
+AutoCodeRover Debug 71.3% 63.2% 30.7% 4.2 28.3 min
+Amazon Q Developer Debug 75.8% 68.4% 49.0% 5.1 35.7 min
+SWE-Agent (GPT-4) Debug 65.2% 57.1% 22.3% 3.8 24.2 min
+Agentic Loop (GPT-4.1) Debug 68.3% 52.1% 17.4% 3.2 22.1 min
+LangGraph + ReAct (GPT-4) Debug 58.7% 45.3% 18.2% 5.4 31.2 min
+LangGraph + ReAct (Claude 4) Debug 61.2% 49.8% 21.3% 6.1 34.5 min
+Kodezi Chronos Debug 89.2% 84.7% 67.3% 7.8 42.3 min
+Key insights from the comprehensive MRR evaluation:
+  1. 1.
+    Debugging vs Code Generation: General-purpose models achieve less than 15% fix accuracy despite having state-of-the-art code generation capabilities.
+This validates our hypothesis that debugging requires fundamentally different architectures.
+  2. 2.
+    LangGraph + ReAct Degradation: Despite being designed for multi-step reasoning, LangGraph+ReAct shows performance degradation in debugging loops (18-21% fix accuracy).
+The ReAct pattern’s observation-thought-action cycle becomes inefficient when debugging requires backtracking and re-evaluation of previous hypotheses.
+  3. 3.
+    Iteration Depth: Chronos performs 7.8 iterations on average compared to 1-2 for general models, demonstrating the importance of persistent debugging loops with execution feedback.
+  4. 4.
+    Precision-Recall Trade-off: While Amazon Q Developer shows impressive 49% fix accuracy, Chronos’s 89.2% precision and 84.7% recall demonstrate superior context identification, leading to 67.3% fix accuracy.
+  5. 5.
+    Time Investment: Chronos takes longer (42.3 min average) but achieves 3-6x higher success rates, making it more time-efficient overall when considering rework costs.
+V-G Adaptive Graph-Guided Retrieval Performance
 We evaluate the impact of AGR on debugging accuracy across different retrieval depths:
-TABLE IV: Performance metrics for different retrieval strategies. Adaptive AGR dynamically selects optimal k based on query complexity.
+TABLE XI: Performance metrics for different retrieval strategies. Adaptive AGR dynamically selects optimal k based on query complexity.
 Retrieval Method k=1 k=2 k=3 k=adaptive Flat
@@
   • Hardware Debugging: Particularly effective for Verilog/VHDL with 91% accuracy (vs 18% baseline)
+111333555101010202020404040505050606060707070808080909090100100100Number of Retrieved NodesPrecision/Recall (%)AGR PrecisionAGR RecallTop-k PrecisionTop-k RecallOracle PrecisionOracle Recall
+Figure 16: Precision-recall trade-off for different retrieval strategies.
+AGR maintains high precision at low k values while achieving comparable recall to oracle retriever, demonstrating efficient noise avoidance.
+V-G1 Why AGR Dominates Existing Graph-Based Retrieval Systems
+While graph-based code retrieval exists in various forms, AGR’s architecture and debugging-specific optimizations make it vastly superior to all existing approaches:
+TABLE XII: AGR vs. existing graph-based retrieval systems on MRR benchmark (5,000 debugging scenarios)
+System Graph Types Debug Success Precision@k=3 Avg Tokens Cross-File Memory Adaptive
+(MRR) (MRR) Retrieved (MRR) Persist Depth
+AGR (Chronos) AST+Log+Test+PR 87.1% 92.8% 31.2K 71.2%
+IntelliCode Compose AST+Import 23.0% 71.0% 62K 18% $\times$ $\times$
+Google Code Search Dependency 23.4% 67.0% 89K 31% $\times$ $\times$
+Facebook Aroma AST 41.0% 78.0% 51K 22% $\times$ $\times$
+DeepMind RETRO Flat vectors 18.0% 52.0% 89K 12% $\times$ $\times$
+MS DeepDebug Stack trace 35.0% 81.0% 21K 28% $\times$ $\times$
+Graph RAG Static KG 28.0% 73.2% 76K 33% $\times$ $\times$
+All results on Multi Random Retrieval (MRR) benchmark with 5,000 real debugging scenarios.
+Debug Success = % of bugs correctly fixed and validated.
+Precision@k=3 = % of retrieved nodes relevant to bug fix.
+Cross-File = % success on bugs spanning multiple files.
+Key Architectural Advantages of AGR:
+1. Unified Multi-Signal Graph Construction
+AGR fuses eight distinct signal types into one traversable graph:
+  • AST parent/child relationships (weight: 0.8)
+  • Import/dependency edges (weight: 0.9)
+  • Function call graphs (weight: 0.85)
+  • Test coverage mappings (weight: 0.95)
+  • Log emission points (weight: 0.92)
+  • Stack trace paths (weight: 0.97)
+  • Commit co-occurrence (weight: $0.7\cdot e^{-\lambda t}$)
+  • PR review comments (weight: 0.6)
+No existing system combines more than 2-3 of these signals.
+This multi-modal fusion enables AGR to trace paths like: test_failure $\rightarrow$ log_entry $\rightarrow$ emitting_function $\rightarrow$ upstream_caller $\rightarrow$ config_error $\rightarrow$ fix_location.
+2. Confidence-Guided Adaptive Expansion
+While other systems use fixed-depth traversal, AGR implements entropy-based adaptive expansion:
+def adaptive_expand(seed_nodes, threshold=0.89):
+"""Expand graph traversal adaptively based on confidence."""
+k = 1
+current_nodes = seed_nodes
+best_nodes_so_far = seed_nodes
+confidence = 0
+while confidence < threshold and k < max_hops:
+neighbors = graph.expand(current_nodes, k)
+confidence = calculate_entropy_confidence(neighbors)
+if confidence > threshold:
+return neighbors
+best_nodes_so_far = neighbors
+current_nodes = neighbors
+k += 1
+return best_nodes_so_far
+This prevents both under-exploration (missing crucial context) and over-exploration (noise flooding). Average expansion: 3.7 hops for complex bugs, 1.2 for simple ones.
+3. Temporal-Aware Edge Weighting
+AGR weights edges based on temporal correlation with debugging success:
+  • Recent bug fixes: Higher weight ($e^{-0.1t}$ decay)
+  • Co-committed files: Boosted during active development
+  • Test-code proximity: Updated after each test run
+  • Historical failure patterns: Accumulated over debugging sessions
+Static systems like IntelliCode or Graph RAG cannot adapt weights based on debugging outcomes.
+4. Persistent Debug Memory Integration
+AGR uniquely interfaces with PDM to:
+  • Cache successful traversal paths (reused in 47ms vs 3.2min initial)
+  • Learn failure-prone subgraphs (87% match rate on recurring bugs)
+  • Accumulate fix patterns (6.8x faster on similar bugs)
+Other systems are stateless, starting from scratch each session.
+5. Debugging-Specific Training and Optimization
+Unlike general-purpose retrievers, AGR is trained on:
+  • 15M real debugging sessions with ground-truth fix locations
+  • Stack trace $\rightarrow$ root cause traversal paths
+  • Test failure $\rightarrow$ code change mappings
+  • Multi-file bug resolution sequences
+This specialization explains the 3-5x performance gap over adapted general retrievers.
+AGR Performance Domination Summary 4x Higher Debug Success: 87.1% vs 23-41% for existing systems 30% Better Precision: 92.8% vs best alternative (81%) 65% Fewer Tokens: 31.2K vs 51-89K average 3x Better Cross-File: 71.2% vs 31% best competitor Only System with Memory: Enables continuous improvement
+V-G2 Oracle Retriever Experiments: Upper Bound Analysis
+To understand the theoretical limits of retrieval-augmented debugging, we conducted experiments with an oracle retriever that has perfect knowledge of which code segments are relevant:
+TABLE XIII: Performance comparison with oracle retriever showing upper bounds for retrieval-based debugging
+Retriever Type Debug Success Precision Recall Avg Context Fix Quality
+Random Baseline 8.7% 31.2% 28.4% 15.2K tokens 2.1/5
+BM25 18.3% 52.8% 49.7% 18.7K tokens 2.8/5
+Dense Retrieval 24.6% 64.3% 61.2% 21.3K tokens 3.2/5
+HyDE 31.2% 71.8% 68.9% 19.8K tokens 3.5/5
+Self-RAG 28.9% 69.4% 66.2% 24.1K tokens 3.4/5
+Graph RAG 33.7% 73.2% 70.8% 26.5K tokens 3.6/5
+AGR (Chronos) 65.3% 89.2% 84.7% 31.2K tokens 4.3/5
+Oracle Retriever 78.9% 100% 100% 42.7K tokens 4.7/5
+Gap to Oracle 13.6% 10.8% 15.3% - 0.4
+Oracle retriever insights:
+  • 21.1% Ceiling: Even with perfect retrieval, 21.1% of bugs cannot be fixed due to reasoning limitations
+  • Chronos Efficiency: AGR achieves 82.7% of oracle performance while using 27% less context
+  • Remaining Gap: The 13.6% gap suggests room for retrieval improvements, particularly in:
+    • Cross-repository dependencies (4.2% of failures)
+    • Implicit behavioral contracts (3.8% of failures)
+    • Domain-specific patterns (2.9% of failures)
+  • Quality vs Quantity: Oracle retrieves more context (42.7K vs 31.2K tokens) but Chronos’s selective retrieval maintains 91% of oracle’s fix quality
+RAG Technique General Tasks Code Tasks Debug Tasks MRR Bench Compute Cost
+Flat Retrieval 71.2% 68.3% 23.4% 31.7% 1.0x
+HyDE 82.1% 74.2% 31.2% 42.3% 2.1x
+Self-RAG 85.7% 78.9% 38.7% 48.1% 1.8x
+FLARE 83.9% 76.5% 35.2% 45.6% 2.3x
+Graph RAG 79.8% 81.2% 41.3% 51.7% 3.2x
+Chronos AGR 88.3% 89.7% 87.1% 89.2% 2.7x
+TABLE XIV: Comparison of RAG techniques across different task types.
+Advanced RAG methods show improvements over flat retrieval, with Chronos’s AGR demonstrating the highest performance on debugging tasks.
 Model HumanEval MBPP Debug Success Root Cause Acc. Retrieval Prec.
-GPT-4 85.7±1.2% 87.0±0.9% 8.5±2.1% 12.3±1.8% 68±2.3%
-GPT-4-Turbo 87.1±1.0% 88.2±0.8% 9.2±1.9% 14.1±1.6% 71±2.0%
-Claude-3-Opus 84.9±1.3% 86.1±1.1% 7.8±2.3% 11.7±2.0% 67±2.4%
-Claude-3-Sonnet 83.2±1.4% 85.2±1.2% 7.3±2.4% 10.9±2.1% 66±2.5%
-Gemini-1.5-Pro 88.3±0.9% 89.1±0.7% 11.2±1.7% 15.8±1.5% 74±1.8%
-CodeT5+ 86.5±1.1% 84.8±1.3% 10.6±1.8% 13.2±1.7% 72±1.9%
-Chronos 90.2±0.6%*** 88.9±0.5%* 65.3±1.4%*** 78.4±1.2%*** 91±0.8%***
-*p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001 compared to best baseline (two-tailed t-test)
-TABLE V: Performance across code synthesis and debugging tasks (mean ± std over 5 runs). Statistical significance shown.
-V-E Comparison with Agentic Code Tools
+GPT-4o 87.8±1.0% 88.5±0.8% 10.3±1.9% 14.7±1.6% 72±2.0%
+GPT-4.1 91.2±0.8% 90.7±0.7% 13.8±1.7% 18.2±1.5% 76±1.8%
+Claude 4 Opus 92.8±0.7% 91.3±0.6% 14.2±1.6% 19.1±1.4% 78±1.7%
+Claude 4 Sonnet 92.1±0.8% 90.9±0.7% 13.6±1.7% 18.5±1.5% 77±1.8%
+DeepSeek V3 90.5±0.9% 89.8±0.8% 12.1±1.8% 16.3±1.6% 75±1.9%
+Qwen2.5-32B 89.7±1.0% 88.9±0.9% 11.5±1.9% 15.7±1.7% 73±2.0%
+Gemini 2.5 Pro 91.6±0.8% 90.2±0.7% 13.9±1.7% 17.9±1.5% 76±1.8%
+Chronos 90.2±0.6%NS 88.9±0.5%NS 65.3±1.4%*** 78.4±1.2%*** 91±0.8%***
+NS: Not significant, *p <0.05<0.05< 0.05, **p <0.01<0.01< 0.01, ***p <0.001<0.001< 0.001 compared to best baseline (two-tailed t-test)
+TABLE XV: Performance across code synthesis and debugging tasks (mean ± std over 5 runs).
+Note that while Chronos shows average performance on pure code generation tasks (HumanEval, MBPP), it dramatically outperforms all models on debugging-specific metrics.
+V-H Comprehensive Debugging Performance Analysis
+To provide a complete picture of debugging capabilities, Table XVI presents detailed metrics across all evaluation dimensions:
+TABLE XVI: Comprehensive debugging performance comparison across all key metrics
+Model/System Root Cause Fix Valid Cross-File MRR Regression Debug Time to
+Precision (%) Accuracy (%) Hit Rate (%) Score Avoid (%) Cycles Fix (min)
+General-Purpose LLMs
+GPT-4.1 31.2±2.1 13.8±1.2 42.3±3.1 0.28 67.2±4.2 8.3±2.1 47.2±12.3
+Claude 4 Opus 34.7±2.3 14.2±1.3 45.8±3.3 0.31 69.8±4.0 7.9±1.9 44.6±11.8
+Gemini 2.5 Pro 29.8±2.0 13.9±1.2 39.2±2.9 0.26 65.3±4.4 9.1±2.3 51.3±13.7
+DeepSeek V3 27.3±1.9 12.1±1.1 37.6±2.8 0.24 63.7±4.5 9.7±2.5 54.8±14.2
+Enhanced RAG Systems
+GPT-4.1 + HyDE 38.9±2.5 22.3±1.8 51.2±3.5 0.36 74.2±3.8 6.2±1.5 38.7±9.4
+Claude 4 + Self-RAG 41.2±2.6 24.7±1.9 53.7±3.6 0.39 76.1±3.6 5.8±1.4 36.2±8.9
+Gemini + GraphRAG 43.8±2.7 28.9±2.1 58.3±3.7 0.42 78.3±3.4 5.3±1.3 33.4±8.1
+Specialized Debugging Tools
+SWE-agent 47.3±2.8 22.9±1.7 61.2±3.8 0.44 81.2±3.2 4.7±1.1 31.2±7.6
+AutoCodeRover 52.1±2.9 31.2±2.2 64.8±3.9 0.48 83.7±3.0 4.2±1.0 28.3±6.9
+LangGraph + ReAct 44.6±2.7 31.2±2.2 56.9±3.7 0.41 79.8±3.3 5.1±1.2 32.6±7.9
+Chronos 87.4±1.2 65.3±1.4 89.2±1.3 0.82 94.6±0.9 2.2±0.4 14.7±3.2
+All metrics averaged over 5,000 debugging scenarios. ± indicates standard error. MRR Score: Multi Random Retrieval benchmark (0-1 scale).
+Key insights from comprehensive analysis:
+  • Root Cause Precision: Chronos achieves 87.4% accuracy in identifying the true source of bugs, 2.5x better than the best baseline
+  • Cross-File Retrieval: 89.2% hit rate demonstrates AGR’s effectiveness at traversing dependencies
+  • Regression Avoidance: 94.6% of Chronos fixes don’t introduce new bugs, critical for production deployment
+  • Efficiency: Average 2.2 debug cycles and 14.7 minutes to fix, showing rapid convergence
+V-I Comparison with Agentic Code Tools
 While traditional LLMs struggle with debugging, a new generation of agentic code tools has emerged. We evaluate Chronos against these systems on real-world debugging scenarios:
-Tool Context Memory Debug Loop Multi-File CI/CD Success Rate
-Cursor Auto IDE only None No Limited No 4.2%
-Claude Code 200K tokens Session No Yes No 6.8%
-OpenAI Codex 8K tokens None No No No 3.1%
-Gemini CLI 1M tokens None No Yes Limited 9.7%
-GitHub Copilot X 16K tokens None No Limited No 5.3%
-Chronos Unlimited* Persistent Yes Yes Yes 65.3%
-TABLE VI: Comparison of Chronos with agentic code tools. *Unlimited via intelligent retrieval and memory.
+TABLE XVII: Comparison of Chronos with modern agentic code tools (2025), sorted by debugging success rate.
+Category Tool Context Memory Debug Loop Multi-File CI/CD Success Rate
+IDE-Integrated Cursor IDE + 32K Session Agent Mode Yes No 4.2%
+Windsurf Cascade Tech Session Write Mode Yes No 5.1%
+Code Assistants GitHub Copilot X 16K tokens None No Limited No 5.3%
+Warp.dev Terminal Session No Limited Yes 7.3%*
+CLI Tools Claude Code CLI 200K tokens Session No Yes No 6.8%
+Gemini CLI 1.5M tokens None No Yes Limited 9.7%
+Debugging-First Chronos Unlimited** Persistent Yes Yes Yes 65.3%
+*Warp achieves 71% on SWE-bench code generation with specialized setup but only 7.3% on real-world debugging tasks (MRR benchmark). **Via AGR intelligent retrieval.
 Key differentiators of Chronos:
@@
   • Unlimited Context: Smart retrieval enables reasoning over entire repositories without token limits
-V-F Cycle-Aware Debugging Case Study
+V-I1 Modern Development Environment Analysis
+The 2025 landscape of AI-powered development tools showcases significant innovation, yet reveals fundamental limitations in debugging capabilities:
+Cursor introduced agent mode for multi-file generation but remains optimized for code completion speed over debugging accuracy.
+Its 4.2% debug success rate reflects prioritization of real-time suggestions over deep reasoning.
+Windsurf by Codeium features Cascade technology for mapping codebases ”like a neural net,” enabling impressive multi-file edits.
+However, its lack of persistent memory and debugging-specific training limits effectiveness on complex bugs.
+Claude Code CLI and Gemini CLI represent the evolution of terminal-based AI assistants.
+Claude Code CLI can use either Opus 4 or Sonnet 4 models, offering flexibility between performance and cost.
+While Claude Code excels at multi-file reasoning and Gemini CLI offers 1M token context, neither implements the iterative debugging loops necessary for autonomous bug resolution.
+Warp.dev demonstrates interesting potential, achieving 71% on SWE-bench with specialized configuration, but this performance doesn’t generalize to real-world debugging scenarios where it achieves only 7.3% success.
+Agentic and Reasoning Models: Despite recent advances, even sophisticated approaches fail at debugging.
+Claude 4 Opus’s agent mode with chain-of-thought reasoning achieves only 16.8% debug success, marginally better than its baseline.
+Reasoning-specific models like o1-preview (18.2%) and o1-mini (15.7%) show similar limitations.
+Web search integration provides modest improvements: Claude with web search reaches 19.3%, while GPT-4.1 with browsing achieves 17.6%.
+These results confirm that debugging requires more than enhanced reasoning or information access, it demands specialized architecture, persistent memory, and iterative refinement capabilities that current models lack.
+V-I2 Comparative Debugging Loop Analysis
+We analyze the debugging approaches of various systems to understand why Chronos achieves superior performance:
+TABLE XVIII: Debugging loop characteristics comparison across systems
+System Loop Type Avg Iter Test Exec Memory Backtrack Success
+Chronos Autonomous 7.8 Persistent 65.3%
+Claude 4 Opus Single-shot 1.2 $\times$ Session $\times$ 14.2%
+GPT-4.1 Limited retry 2.1 $\times$ Session $\times$ 13.8%
+Cursor User-driven 3.5 Manual Session $\times$ 4.2%
+Windsurf Multi-file 2.8 Manual Session $\times$ 6.1%
+Amazon Q Guided 4.2 Session Limited 49.0%
+GitHub Copilot Suggestion 1.0 $\times$ None $\times$ 8.7%
+Warp.dev Config-based 3.1 Session $\times$ 7.3%
+Key observations:
+  • Iteration Depth: Chronos averages 7.8 iterations vs 1-4 for others, enabling deeper exploration
+  • Test Integration: Only Chronos, Amazon Q, and Warp.dev execute tests automatically
+  • Backtracking: Chronos uniquely supports full hypothesis backtracking when fixes fail
+  • Memory Persistence: All competitors use session-based memory, losing context between runs
+V-J Quantitative Analysis: Debug Cycles and Convergence Rates
 To showcase Chronos’s real-world debugging prowess, we conduct a qualitative study involving a set of regression bug scenarios drawn from a large open-source Python project. Metrics include:
@@
   • Time-to-resolution and reduction in manual engineering effort.
-GPT-4ClaudeCodeT5+Chronos0022224444666688886.56.56.56.56.86.86.86.85.15.15.15.12.22.22.22.2ModelAvg. Cycles to Fix
+GPT-4.1Claude 4 OpusDeepSeek V3Chronos02224446668884.84.84.84.54.54.54.24.24.22.22.22.2ModelAvg. Cycles to Fix
 Figure 17: Average code-to-fix cycles for Chronos and baseline models on real-world bugs (lower is better).
-Chronos demonstrates not only higher accuracy and retrieval precision, but also a dramatically reduced number of debug cycles—underscoring its benefit for continuous, automated codebase reliability.
-V-G Case Studies
+Chronos demonstrates not only higher accuracy and retrieval precision, but also a dramatically reduced number of debug cycles, underscoring its benefit for continuous, automated codebase reliability.
+V-K Real-World Debugging Scenarios: Detailed Case Analysis
 To illustrate Chronos’s debugging capabilities, we present two real-world examples from our evaluation:
-V-G1 Case Study 1: Cross-Module Null Pointer Exception
+V-K1 Case Study 1: Cross-Module Null Pointer Exception
 Bug Report: ”Application crashes with NullPointerException when processing user exports after recent authentication refactor”
+Technical Analysis:
+  • Initial Context: 3,247 tokens from stack trace, 2 affected files
+  • AGR Retrieval: k=1 found 3 auth commits, k=2 expanded to 7 related files, k=3 included test cases
+  • Confidence Progression: 32% $\rightarrow$ 67% $\rightarrow$ 91% (terminated at k=3)
 Chronos Reasoning Process:
   1. 1.
-    Retrieved recent commits mentioning ”authentication” (found 3 relevant commits)
+    Semantic Search (142ms): Retrieved commits with ”authentication” keyword via Memory Engine
   2. 2.
-    Traced data flow from auth module to export module
+    Data Flow Analysis (287ms): Traced AuthToken propagation through 5 call sites
   3. 3.
-    Identified missing null check after auth token refresh
+    Pattern Recognition (95ms): Identified missing null-safety pattern post-refactor
   4. 4.
-    Located similar pattern in 2 other modules
+    Impact Analysis (178ms): Located 2 similar vulnerabilities in UserService and AdminService
 Generated Fix:
-Null Pointer Fix in Authentication Flow // In AuthService.java public AuthToken refreshToken(String userId) { AuthToken token = tokenCache.get(userId); if (token == null $|$ token.isExpired()) { token = authProvider.generateToken(userId); + if (token != null) { + tokenCache.put(userId, token); + } } return token; } // In ExportService.java public void exportUserData(String userId) { AuthToken token = authService.refreshToken(userId); + if (token == null) { + throw new AuthenticationException( + "Failed to authenticate user for export"); + } // ...
-rest of export logic }
-Validation Results: All 47 test cases passed, including 3 new tests Chronos generated for the edge case.
-V-G2 Case Study 2: Async Race Condition in Message Queue
+Null Pointer Fix in Authentication Flow
+// In AuthService.java
+public AuthToken refreshToken(String userId) {
+AuthToken token = tokenCache.get(userId);
+if (token == null || token.isExpired()) {
+token = authProvider.generateToken(userId);
++ if (token != null) {
++ tokenCache.put(userId, token);
+return token;
+// In ExportService.java
+public void exportUserData(String userId) {
+AuthToken token = authService.refreshToken(userId);
++ if (token == null) {
++ throw new AuthenticationException(
++ "Failed to authenticate user for export");
+// ... rest of export logic
+Validation Results:
+  • Test Coverage: 47/47 existing tests passed, 3 new edge case tests generated
+  • Performance Impact: No measurable latency increase (<1<1< 1ms overhead)
+  • Fix Iterations: Converged in 2 cycles (initial fix $\rightarrow$ test generation $\rightarrow$ validation)
+  • Total Time: 1.7 seconds from bug report to validated fix
+V-K2 Case Study 2: Async Race Condition in Message Queue
 Bug Report: ”Intermittent message loss in high-load scenarios, approximately 0.1% messages not processed”
-Chronos Analysis:
-  • Retrieved message queue implementation and recent load test results
-  • Identified race condition between acknowledgment and connection pooling
-  • Found similar issue fixed 8 months ago in different component
-  • Applied analogous solution with adaptations
+Technical Analysis:
+  • Initial Context: 8,432 tokens (logs, metrics, queue implementation)
+  • AGR Retrieval: k=4 required due to distributed system complexity
+  • Pattern Matching: Found similar fix in commit a3f8b2c via Memory Engine
+Chronos Debugging Process:
+  1. 1.
+    Log Analysis (523ms): Correlated message IDs with acknowledgment timestamps
+  2. 2.
+    Concurrency Detection (1,247ms): Identified non-atomic operation between ack() and releaseConnection()
+  3. 3.
+    Historical Search (432ms): Retrieved 3 similar race conditions from Memory Engine
+  4. 4.
+    Solution Synthesis (298ms): Adapted previous synchronization pattern to current architecture
 Generated Fix:
-Race Condition Fix in Message Queue // MessageProcessor.java private void processMessage(Message msg) { try { handler.process(msg); - connectionPool.returnConnection(conn); - msg.acknowledge(); + // Fix: Acknowledge before returning connection + msg.acknowledge(); + connectionPool.returnConnection(conn); } catch (Exception e) { + // Ensure connection returned even on error + connectionPool.returnConnection(conn); msg.nack(); throw e; } }
+Race Condition Fix in Message Queue
+// MessageProcessor.java
+private void processMessage(Message msg) {
+try {
+handler.process(msg);
+- connectionPool.returnConnection(conn);
+- msg.acknowledge();
++ // Fix: Acknowledge before returning connection
++ msg.acknowledge();
++ connectionPool.returnConnection(conn);
+} catch (Exception e) {
++ // Ensure connection returned even on error
++ connectionPool.returnConnection(conn);
+msg.nack();
+throw e;
 Results: Load test with 10M messages showed 0
-V-H Long-Context Debugging Performance
+V-K3 Complete Debugging Trace: Memory Persistence Example
+To demonstrate Chronos’s full debugging workflow and memory retention, we present a complete trace from our evaluation:
+[…]
+PDM Match: 87% similarity to ”Redis timestamp precision”
+Chronos: Retrieved previous fix, adapted to session context
+Time to Fix: 47 seconds (vs 3.2 minutes for original)
+This trace demonstrates how PDM enables Chronos to learn from past debugging sessions, dramatically reducing fix time for similar issues.
+V-L Performance at Scale: Million-Token Context Evaluation
 Even models with extended context windows fail at real debugging tasks due to fundamental architectural limitations:
 Model Context Size Cross-File Bugs Historical Bugs Complex Traces Avg. Success
-GPT-4-32K 32K tokens 7.2±1.9% 3.1±1.2% 5.8±1.6% 5.4±1.5%
-Claude-3-200K 200K tokens 9.8±2.1% 4.7±1.4% 8.3±1.8% 7.6±1.7%
-Gemini-1.5-Pro-1M 1M tokens 14.3±2.3% 6.2±1.5% 11.7±2.0% 10.7±1.9%
-Chronos Unlimited 71.2±1.8%*** 68.9±2.0%*** 74.3±1.6%*** 71.5±1.8%***
-***p ¡ 0.001 compared to Gemini-1.5-Pro-1M (paired t-test, n=100 tasks per category)
-TABLE VII: Performance on debugging tasks requiring extensive context. Statistical significance demonstrates retrieval superiority.
+GPT-4.1 128K tokens 15.7±2.2% 7.8±1.6% 13.2±1.9% 12.2±1.8%
+Claude 4 Opus 200K tokens 16.2±2.1% 8.3±1.5% 14.1±1.9% 12.9±1.8%
+DeepSeek V3 128K tokens 13.5±2.2% 6.9±1.6% 11.8±2.0% 10.7±1.9%
+Gemini 2.5 Pro 1M tokens 17.1±2.1% 9.2±1.5% 15.3±1.8% 13.9±1.7%
+Chronos Unlimited* 71.2±1.8%*** 68.9±2.0%*** 74.3±1.6%*** 71.5±1.8%***
+*Via AGR. ***p <0.001<0.001< 0.001 compared to Gemini 2.5 Pro (paired t-test, n=100 tasks per category)
+TABLE XIX: Performance on debugging tasks requiring extensive context.
+Despite models having up to 1M tokens, intelligent retrieval and debug-specific training enable Chronos to achieve 5x better performance.
 The results demonstrate that raw context size alone cannot solve debugging.
-Chronos’s intelligent retrieval, persistent memory, and debug-specific training enable it to outperform even million-token models by over 6x.
-V-I Detailed Performance Analysis
+Chronos’s intelligent retrieval, persistent memory, and debug-specific training enable it to outperform even million-token models by over 5x.
+V-M Detailed Performance Analysis
 We further analyze Chronos’s performance across different bug categories and complexity levels:
 Bug Category Syntax Logic Concurrency Memory API Performance
-GPT-4 82.3% 12.1% 3.2% 5.7% 18.9% 7.4%
-Claude-3-Opus 79.8% 10.7% 2.8% 4.3% 16.2% 6.1%
-Gemini-1.5-Pro 85.1% 15.3% 4.1% 6.9% 22.4% 9.8%
+GPT-4.1 88.7% 17.3% 5.8% 8.2% 24.6% 11.3%
+Claude 4 Opus 91.2% 18.9% 6.3% 9.1% 26.8% 12.7%
+DeepSeek V3 87.9% 16.2% 4.9% 7.6% 23.1% 10.5%
+Qwen2.5-32B 86.8% 15.7% 4.5% 7.2% 22.3% 9.8%
+Gemini 2.5 Pro 89.3% 17.8% 5.7% 8.7% 25.3% 11.9%
 Chronos 94.2% 72.8% 58.3% 61.7% 79.1% 65.4%
-TABLE VIII: Success rates by bug category. Chronos shows particular strength in complex bug types that require deep understanding.
-Repository Size ¡10K LOC 10K-100K 100K-1M ¿1M LOC
-GPT-4 15.2% 9.8% 4.3% 1.2%
-Claude-3-200K 17.8% 11.2% 5.7% 2.1%
-Gemini-1.5-Pro 21.3% 14.7% 8.9% 3.8%
+TABLE XX: Success rates by bug category.
+While frontier models show incremental improvements over their predecessors, Chronos demonstrates 3-10x better performance on complex bug types through debug-specific training.
+[…]
+GPT-4.1 18.5% 13.2% 7.8% 3.1%
+Claude 4 Opus 21.7% 15.8% 9.2% 4.3%
+DeepSeek V3 19.2% 13.9% 8.1% 3.5%
+Gemini 2.5 Pro 24.1% 17.3% 11.5% 5.2%
 Chronos 71.2% 68.9% 64.3% 59.7%
-TABLE IX: Debugging success rates by repository size, demonstrating Chronos’s scalability.
-V-J Multi-Code Association Retrieval Performance
+TABLE XXI: Debugging success rates by repository size, demonstrating Chronos’s scalability.
+V-N Multi-Code Association Retrieval Performance
 We evaluate Chronos’s ability to retrieve and associate multiple code artifacts for debugging:
@@
 Average 90.5±0.8% 90.1±0.9% 90.3±0.7%
-TABLE X: Multi-code association retrieval performance across different debugging contexts.
-V-K Computational Efficiency
+TABLE XXII: Multi-code association retrieval performance across different debugging contexts.
+V-O Efficiency Metrics: Cost, Latency, and Resource Usage
 A critical consideration for production deployment is computational efficiency. We analyze Chronos’s performance characteristics compared to baselines and human debugging:
-Metric GPT-4 Claude-3 Gemini-1.5 Chronos Human Dev
-Avg. Time to Fix 82.3s 76.9s 71.2s 134.7s 2.4 hours
+Metric GPT-4.1 Claude 4 Opus Gemini 2.5 Pro Chronos Human Dev
+Avg. Time to Fix 68.5s 64.2s 59.8s 134.7s 2.4 hours
 Context Window 128K tokens 200K tokens 1M tokens Unlimited* N/A
-Cost per Bug \$0.47 \$0.52 \$0.68 \$0.89 \$180
-Success Rate 8.5% 7.8% 11.2% 65.3% 94.2%
-Effective Cost* \$5.53 \$6.67 \$6.07 \$1.36 \$191
+Cost per Bug \$0.52 \$0.58 \$0.72 \$0.89 \$180
+Success Rate 13.8% 14.2% 13.9% 65.3% 94.2%
+Effective Cost* \$3.77 \$4.08 \$5.18 \$1.36 \$191
 *Unlimited via dynamic retrieval; Effective cost = Cost per bug / Success rate
-TABLE XI: Computational efficiency and cost analysis. Despite higher per-attempt cost, Chronos’s high success rate yields lowest effective cost.
-V-K1 Inference Time Breakdown
+TABLE XXIII: Computational efficiency and cost analysis. Despite higher per-attempt cost, Chronos’s high success rate yields lowest effective cost.
+V-O1 Inference Time Breakdown
 Chronos’s 134.7s average debugging time consists of:
@@
   • Memory Update: 12.3s (9.1%)
-V-K2 Return on Investment Analysis
+V-O2 Return on Investment Analysis
 For a typical enterprise with 100 developers:
@@
   • ROI: 47:1 in first year, accounting for infrastructure and licensing
-VI DISCUSSION AND ABLATION ANALYSIS
-VI-A Qualitative Insights
+V-O3 Detailed Inference Cost and Latency Analysis
+We conducted comprehensive benchmarking of inference costs and latency across different bug complexity levels:
+TABLE XXIV: Inference cost and latency breakdown by bug complexity and system
+System Simple Bugs Medium Bugs Complex Bugs
+Cost (\$) Latency (s) Cost (\$) Latency (s) Cost (\$) Latency (s)
+GPT-4.1 (128K) 0.18 12.3 0.52 68.5 1.24 187.2
+Claude 4 Opus 0.21 14.7 0.58 64.2 1.41 201.3
+Gemini 2.0 Pro 0.31 8.9 0.72 59.8 1.93 156.4
+Amazon Q Dev 0.42 21.3 0.91 94.7 2.14 234.6
+Chronos (Ours) 0.34 31.2 0.89 134.7 1.78 298.4
+- Retrieval only 0.08 7.8 0.19 23.4 0.41 52.3
+- LLM inference 0.21 18.9 0.56 67.8 1.12 156.8
+- Test execution 0.03 3.1 0.09 31.2 0.18 67.2
+- Memory update 0.02 1.4 0.05 12.3 0.07 22.1
+Cost-Effectiveness Analysis (Cost / Success Rate)
+GPT-4.1 1.30 - 3.77 - 8.99 -
+Claude 4 Opus 1.48 - 4.08 - 9.93 -
+Gemini 2.0 Pro 2.23 - 5.18 - 13.88 -
+Amazon Q Dev 0.86 - 1.86 - 4.37 -
+Chronos 0.52 - 1.36 - 2.73 -
+Key observations:
+  • Latency vs Accuracy Trade-off: Chronos’s higher latency (2-3x) is offset by 4-5x better success rates
+  • Cost Scaling: Complex bugs show sublinear cost increase (5.2x) despite exponential difficulty
+  • Component Distribution: LLM inference dominates cost (63%) while retrieval dominates initial latency
+  • Parallelization Opportunity: Test execution (23% of latency) can be parallelized for 1.3x speedup
+VI ANALYSIS: WHY CHRONOS SUCCEEDS WHERE OTHERS FAIL
+VI-A The Debugging Specialization Hypothesis: Evidence and Implications
+The stark performance gap between Chronos (65.3% debug success) and state-of-the-art general models ($\leq$15%) reveals a fundamental truth: debugging is not simply an extension of code generation.
+While models like Claude 4 Opus (72.5% on SWE-bench) and GPT-4.1 (54.6%) excel at writing new code, debugging demands distinct capabilities:
+  1. 1.
+    Temporal Reasoning: Understanding how code evolved through commits, why certain patterns were introduced, and which changes correlate with bug emergence.
+General models lack this historical perspective.
+  2. 2.
+    Multi-Modal Signal Integration: Debugging requires synthesizing error traces, logs, test failures, and code, a fundamentally different task than generating syntactically correct code from specifications.
+  3. 3.
+[…]
+  4. 4.
+[…]
+VI-B Key Architectural Insights: Memory, Iteration, and Context
 Our evaluation highlights several domains where Chronos delivers outsized impact compared to prior systems:
@@
   • Continuous Knowledge Incorporation: By feeding CI/CD, reviewer, and test feedback into persistent memory, Chronos improves its project-specific performance over time, exhibiting lower repeated error rates and faster adaptation to new code patterns.
-Bug Scenario GPT-4 Chronos Chronos Resolution Path
+Bug Scenario GPT-4.1 Chronos Chronos Resolution Path
 Test Failure on user_auth Incorrect var patch Full fix Traced import drift $\rightarrow$ found stale config $\rightarrow$ auto-fix and doc update
@@
 Intermittent CI Error Flaky retry logic Full fix Ingested CI logs, patched async boundary, added test case and explanation
-TABLE XII: Qualitative examples where Chronos successfully applies multi-code context to resolve debugging tasks beyond the reach of baseline LLMs.
-VI-B Ablation Studies
+TABLE XXV: Qualitative examples where Chronos successfully applies multi-code context to resolve debugging tasks beyond the reach of baseline LLMs.
+VI-C Ablation Studies
 To isolate the contribution of core design features, we perform targeted ablations:
@@
   • No Orchestration Loop: Disabling the validate-retrieve-update workflow reverts performance to basic code suggestion with higher error rate and longer time-to-fix.
-FullNo MCAStaticNo Orch002020202040404040606060608080808010010010010090909090494949496262626255555555ConfigurationDebug Success (%)
+TABLE XXVI: Ablation study showing contribution of each component. ✓ indicates component enabled.
+Model Variant Memory AGR Test Loop Fix Rate↑ Bug Loc.↑
+Full Chronos 65.3% 87.2%
+w/o Memory $\times$ 48.1% (-17.2) 71.5% (-15.7)
+w/o AGR $\times$ 42.6% (-22.7) 63.8% (-23.4)
+w/o Test Loop $\times$ 51.7% (-13.6) 79.3% (-7.9)
+w/o Memory+AGR $\times$ $\times$ 31.2% (-34.1) 52.1% (-35.1)
+w/o Memory+Test $\times$ $\times$ 35.8% (-29.5) 58.7% (-28.5)
+w/o AGR+Test $\times$ $\times$ 33.4% (-31.9) 55.2% (-32.0)
+Baseline (None) $\times$ $\times$ $\times$ 22.1% (-43.2) 41.3% (-45.9)
+FullNo MCAStaticNo Orch0202020404040606060808080100100100909090494949626262555555ConfigurationDebug Success (%)
 Figure 18: Ablation analysis: Debugging success rate with each Chronos core component removed (lower is worse).
 These findings underscore the essential synergy between deep memory, multi-code contextualization, and autonomous workflow orchestration for effective debugging and adaptive code maintenance.
-VI-C Failure Analysis
+VI-D Failure Mode Analysis: Understanding Chronos’s Limitations
 Despite Chronos’s strong performance, our analysis reveals specific failure modes and bug categories where the system struggles:
-VI-C1 Common Failure Modes
+VI-D1 Common Failure Modes
   1. 1.
@@
     Issue: Chronos lacks healthcare regulatory knowledge
-VI-C2 Edge Cases and Limitations
-  • Extremely Large Monorepos (¿10M LOC): Performance degrades to 45.3% success rate due to retrieval precision issues
+VI-D2 Edge Cases and Limitations
+  • Extremely Large Monorepos ($>10M$ LOC): Performance degrades to 45.3% success rate due to retrieval precision issues
   • Legacy Code with Poor Documentation: Success drops to 38.9% when code lacks comments and uses cryptic variable names
@@
   • UI/UX Bugs: Visual rendering issues essentially unsolvable (8.3% success) without screenshot analysis
+Concrete Failure Case Study: Chronos struggled with a TypeScript monorepo using Lerna, dynamic imports, and cross-package configuration resolution.
+Despite correctly retrieving all relevant files across packages, PDM lacked adequate context to resolve type scope ambiguity when multiple packages exported identically-named interfaces.
+The fix attempted to import from the wrong package, causing circular dependencies.
+This failure mode highlights the need for semantic-aware static linking and better understanding of module resolution strategies in complex build systems.
+Future work should incorporate build tool semantics directly into the memory layer.
 Bug Category Success Rate Primary Failure Reason
@@
 UI/Visual 8.3±1.9% No visual understanding
-TABLE XIII: Chronos performance on challenging bug categories.
-VII LIMITATIONS, FUTURE WORK, AND BROADER IMPACT
-VII-A Current Limitations
-While Chronos marks a significant advance in autonomous code maintenance, certain limitations remain:
-  • Extreme-Scale Context Latency: Despite highly efficient memory and retrieval, performance can be constrained in ultra-large monolithic repositories during peak concurrent update hours.
-  • Initial Memory Cold Start: Brand new projects (no history) or projects with very sparse documentation may exhibit reduced multi-code association richness, necessitating a brief period of active learning.
-  • Non-Determinism and Overcorrection: In some edge cases, especially with noisy historical feedback, Chronos has shown a tendency to ”over-correct” or introduce minor regressions before self-stabilization.
-  • Opaque Reasoning Pathways: The fully autonomous nature and integration of deep learning sometimes hinder model explainability, which may limit adoption in highly regulated or safety-critical industries.
-VII-B Future Work
-We are actively addressing these limitations and have several concrete directions for improvement:
-  • Incremental Embedding Optimization: Research into hybrid memory systems and smarter incremental re-indexing to further minimize retrieval and context assembly latency during scale-out operations.
-  • Interactive and Visual Explanations: Implementation of stepwise reasoning visualizers and debug trace replays, making Chronos’ decision process more transparent to engineers and auditors.
-  • Multi-Project/Organization Federated Memory: Extending Chronos memory to operate across multiple repositories and organizational boundaries, supporting federated bug detection, code migration, and systemic risk profiling.
-  • Human-in-the-Loop Collaboration: Exploring adaptive workflows where engineers and Chronos can jointly control the loop—overriding, annotating, or confirming automated fixes in sensitive code regions.
-  • Security and Adversarial Robustness: Building rigorous defense mechanisms against prompt injection, memory poisoning, and adversarial debugging cases.
-VII-C Deployment Architecture and Integration
+TABLE XXVII: Chronos performance on challenging bug categories.
+VI-E Adversarial Evaluation
+To assess Chronos’s robustness against malicious inputs and poisoned training data, we conduct comprehensive adversarial testing:
+VI-E1 Attack Methodology
+We evaluate three categories of adversarial attacks:
+1. Input Perturbation Attacks:
+  • Prompt Injection: Inserting malicious instructions in bug descriptions
+  • Code Obfuscation: Deliberately confusing variable names and control flow
+  • Misleading Comments: Comments that contradict actual code behavior
+2. Data Poisoning Attacks:
+  • Backdoor Triggers: Specific code patterns that trigger incorrect fixes
+  • Label Flipping: Training examples with intentionally wrong fixes
+  • Gradient Attacks: Adversarial examples crafted to maximize loss
+3. Retrieval Manipulation:
+  • Context Stuffing: Flooding retrieval with irrelevant but similar code
+  • Dependency Confusion: Creating fake dependencies to mislead AGR
+  • Temporal Attacks: Manipulating commit timestamps to affect retrieval
+VI-E2 Robustness Results
+TABLE XXVIII: Adversarial robustness evaluation on 1,000 attack samples per category
+Category Attack Type Success Detection Mitigation
+Rate Rate
+Input Prompt Injection 12.3% 87.7% Input sanitization
+Perturbation Code Obfuscation 23.1% 76.9% AST normalization
+Misleading Comments 8.7% 91.3% Code-comment align
+Data Backdoor Triggers 5.2% 94.8% Anomaly detection
+Poisoning Label Flipping 15.6% 84.4% Consistency checks
+Gradient Attacks 19.3% 80.7% Gradient clipping
+Retrieval Context Stuffing 31.4% 68.6% Relevance filtering
+Manipulation Dependency Confusion 27.8% 72.2% Graph validation
+Temporal Attacks 11.2% 88.8% Timestamp verify
+Key Findings:
+  • AGR’s graph structure provides natural defense against context stuffing (68.6% detection)
+  • PDM’s pattern matching detects 94.8% of backdoor triggers through anomaly scores
+  • Temporal attacks are largely ineffective due to multi-signal validation
+VI-F Scalability Analysis
+We evaluate Chronos’s performance across codebases ranging from 1K to 100M lines of code:
+VI-F1 Performance Metrics by Scale
+TABLE XXIX: Scalability analysis across different codebase sizes
+Codebase Size Debug Avg Response Memory Graph Build Index
+(LOC) Success Time (s) Usage (GB) Time (min) Size (GB)
+1K-10K 91.2% 2.3 0.5 0.1 0.01
+10K-100K 89.7% 4.7 2.1 1.2 0.08
+100K-1M 87.1% 8.9 8.7 12.4 0.92
+1M-10M 82.3% 18.2 31.5 87.3 9.7
+10M-100M 73.4% 45.7 128.3 512.8 98.2
+VI-F2 Optimization Strategies at Scale
+  1. 1.
+    Incremental Graph Updates: For codebases $>1M$ LOC, we implement differential graph updates that process only changed files, reducing update time by 87%.
+  2. 2.
+    Hierarchical Indexing: Multi-level indexes with module-level summaries enable sub-linear retrieval complexity O(log n) for repositories $>10M$ LOC.
+  3. 3.
+    Distributed Processing: Graph construction parallelizes across 32 cores, achieving near-linear speedup for codebases up to 50M LOC.
+  4. 4.
+    Memory-Mapped Storage: PDM uses memory-mapped files for patterns exceeding RAM, maintaining <100<100< 100ms access latency.
+VI-F3 Detailed Memory Consumption Analysis
+We profile memory usage across different components and operations:
+TABLE XXX: Memory consumption breakdown by component (1M LOC codebase)
+Component Peak Memory % of Total Growth Rate
+AGR Graph Structure 2.8 32.2% O(n log n)
+PDM Pattern Storage 1.9 21.8% O(m)
+Embedding Cache 1.6 18.4% O(n)
+AST Representations 1.2 13.8% O(n)
+Temporal Indexes 0.7 8.0% O(n log t)
+Runtime Buffers 0.5 5.8% O(1)
+Total 8.7 100% -
+Memory Optimization Strategies:
+  • Lazy Loading: Graph nodes loaded on-demand, reducing baseline memory by 65%
+  • Pattern Compression: PDM patterns compressed using AST-aware encoding (3.2x compression ratio)
+  • Cache Eviction: LRU policy for embeddings with 90% hit rate at 20% memory cost
+  • Incremental GC: Custom garbage collection for graph traversal reduces peak memory by 40%
+VI-G Human Evaluation Study
+We conducted a controlled study with N=50 professional developers to assess Chronos’s real-world effectiveness:
+VI-G1 Study Design
+Participants: 50 developers (5-15 years experience) from 12 companies
+  • 20 backend engineers (Java/Python)
+  • 15 full-stack developers (JavaScript/TypeScript)
+  • 10 infrastructure engineers (Go/Rust)
+  • 5 ML engineers (Python/C++)
+Tasks: Each developer debugged 10 real production bugs:
+  • 5 bugs with Chronos assistance
+  • 5 bugs with their preferred tools (baseline)
+  • Randomized order to prevent learning effects
+  • Bugs selected from actual production incidents
+VI-G2 Quantitative Results
+TABLE XXXI: Human evaluation results (N=50 developers, 500 debugging sessions)
+Metric With Chronos Baseline Improvement p-value
+Fix Success Rate 87.2% 62.4% +39.7% <0.001<0.001< 0.001
+Time to Fix (min) 18.3±7.2 43.7±19.8 -58.1% <0.001<0.001< 0.001
+Code Quality Score 8.7/10 7.2/10 +20.8% <0.01<0.01< 0.01
+Confidence Rating 8.9/10 6.8/10 +30.9% <0.001<0.001< 0.001
+Would Use Again 92% - - -
+VI-G3 Qualitative Feedback
+Positive Themes:
+  • ”AGR found cross-file dependencies I would have missed” (31/50 developers)
+  • ”PDM patterns saved hours on recurring issues” (27/50)
+  • ”Explanations helped me understand the fix” (42/50)
+Areas for Improvement:
+  • ”Needs better IDE integration” (18/50)
+  • ”Sometimes retrieves too much context” (12/50)
+  • ”UI lag on very large codebases” (8/50)
+VI-H Fine-grained Ablations
+We conduct detailed ablations on individual AGR components to understand their contributions:
+VI-H1 Component-wise Ablation Results
+TABLE XXXII: Fine-grained ablation study on AGR components (MRR benchmark)
+Category Configuration Debug Prec@3 Hops
+Baseline Full AGR System 87.1% 92.8% 2.3
+Graph Remove AST edges 71.2% 84.1% 3.1
+Construction Remove log flow edges 78.9% 87.3% 2.7
+Remove test trace edges 73.4% 81.2% 3.8
+Remove temporal weights 82.3% 89.7% 2.5
+Traversal Fixed k=3 (no adaptive) 69.8% 78.4% 3.0
+Strategy No confidence weighting 74.1% 83.2% 4.2
+BFS instead of guided 61.3% 71.8% 5.7
+No backtracking 79.2% 88.1% 2.8
+Memory No PDM patterns 72.8% 85.3% 3.2
+Integration No pattern ranking 81.2% 90.1% 2.6
+No temporal decay 83.7% 91.2% 2.4
+Critical Findings:
+  1. 1.
+    AST edges are most critical: Removing them causes 18.3% performance drop
+  2. 2.
+    Adaptive depth is essential: Fixed k=3 reduces success by 19.9%
+  3. 3.
+    PDM integration provides 16.4% boost: Pattern memory significantly improves debugging
+  4. 4.
+    Guided traversal beats BFS by 29.6%: Intelligence in path selection matters
+VII DESIGN RATIONALE AND THEORETICAL MOTIVATION FOR AGR
+The Adaptive Graph-Guided Retrieval (AGR) system addresses three fundamental limitations of vector-based retrieval in debugging contexts.
+First, real codebases exhibit structural dependencies that span multiple files, a bug in AuthService.java may stem from a configuration change in config.yml that affects DatabasePool.java, which then impacts authentication.
+Traditional top-k vector search fails here because these files share minimal textual similarity.
+AGR’s multi-hop traversal follows actual code dependencies (imports, function calls, data flow), recovering the complete causal chain regardless of embedding distances.
+Second, local neighborhood propagation in AGR exploits the insight that code proximity (in the dependency graph) correlates strongly with debugging relevance.
+When a test fails in PaymentTest.java, the bug is exponentially more likely to be within 1-3 hops in the dependency graph than in a random file with high textual similarity.
+This locality principle, combined with confidence-weighted traversal, enables AGR to achieve 89% precision where vector search achieves only 31%.
+Third, temporal anchoring prevents the retrieval of stale code patterns in actively maintained repositories.
+By weighting edges based on commit recency and co-modification frequency, AGR naturally prioritizes current architectural patterns over deprecated ones, reducing false positives by 67% in repositories with >1000>1000> 1000 commits/month.
+Bug: NPE in ExportExportService.javaExportUtils.javaExportConfig.java×$\times$ Misses Authsim=0.89sim=0.72sim=0.68Vector RAG: Top-3 SimilarBug: NPE in ExportExportService.javaAuthService.javaTokenCache.java✓ Found Root Causehop=0callsusesrecentcommitAGR: Dependency Traversal3x better
+Figure 19: AGR vs Vector RAG: While vector search retrieves textually similar files that miss the root cause, AGR follows actual code dependencies and temporal signals to locate the authentication bug causing the export failure.
+The superiority of AGR demonstrated in Figure 19 is not merely empirical, it stems from fundamental theoretical properties that we now formalize.
+VIII THEORETICAL GUARANTEES
+Having demonstrated AGR’s practical superiority through extensive evaluation, we now provide theoretical foundations that explain why these improvements are fundamental rather than incidental.
+VIII-A Why AGR Converges Efficiently
+[…]
+Convergence Guarantee: Given a code graph with maximum degree $d$, AGR converges to the optimal debugging context within $k$ iterations (typically k=5) with probability exceeding 98%.
+Practical Implications:
+  • Bounded search: AGR examines at most $O(d^{k})$ nodes, preventing exponential blowup
+  • Early termination: 73% of bugs are resolved within 3 hops
+  • Predictable performance: Response time scales logarithmically with codebase size
+VIII-B Fix Correctness with PDM
+The Persistent Debug Memory provides statistical guarantees on fix quality:
+Pattern Matching Accuracy: With $m$ stored bug patterns, the probability of generating a correct fix increases as:
+$P(\text{correct fix})\geq 1-e^{-m/1000}$
+[…]
+VIII-C Why This Matters for Production Debugging
+Unlike theoretical debugging models, AGR’s guarantees translate directly to production benefits:
+TABLE XXXIII: Theoretical guarantees vs. real-world impact
+Theoretical Property Guarantee Production Impact
+Convergence time O(k log d) <10<10< 10s for 1M LOC
+[…]
+Memory usage O(n log n) Fits in 32GB RAM
+[…]
+IX FAILURE ANALYSIS: DEEP DIVE INTO CHALLENGING CASES
+While our theoretical guarantees and empirical results demonstrate Chronos’s effectiveness, understanding failure modes is crucial for both users and future improvements.
+We conduct systematic failure analysis on 2,500 debugging sessions where Chronos failed to produce correct fixes:
+IX-A Failure Taxonomy
+TABLE XXXIV: Detailed failure analysis across 2,500 failed debugging sessions
+Failure Category Subcategory Count % of Failures Root Cause
+Retrieval Failures
+(42.3% of total)
+Missing Context 423 16.9% AGR depth limit reached
+Wrong Context 312 12.5% Semantic aliasing confusion
+Stale Context 323 12.9% Outdated cached patterns
+Understanding Failures
+(31.2% of total)
+Complex Logic 287 11.5% Multi-step reasoning limit
+Domain Specific 234 9.4% Missing specialized knowledge
+Implicit Behavior 256 10.2% Undocumented assumptions
+Generation Failures
+(26.5% of total)
+Incorrect Fix 298 11.9% Wrong root cause identified
+Partial Fix 189 7.6% Incomplete solution
+Breaking Changes 176 7.0% Side effects not considered
+IX-B Representative Failure Cases
+IX-B1 Case 1: Distributed Consensus Bug
+Bug Description: Raft consensus implementation fails during leader election with 5+ nodes
+Failure Mode: Chronos retrieved all relevant consensus code but failed to identify the subtle race condition in vote counting across network partitions.
+// Chronos attempted fix (incorrect)
+if (voteCount > clusterSize / 2) {
+becomeLeader(); // Missing: partition check
+// Correct fix required understanding of:
+// 1. Network partition detection
+// 2. Split-brain prevention
+// 3. Distributed state consistency
+Listing 1: Distributed consensus bug: Chronos’s incorrect fix
+Root Cause: Lacks theoretical understanding of distributed systems safety properties and FLP impossibility theorem implications.
+IX-B2 Case 2: Memory Corruption in C++ Template Metaprogramming
+Bug Description: Segfault in recursive template instantiation with variadic parameters
+Failure Mode: AGR correctly identified template definitions but PDM had no patterns for template instantiation depth issues.
+// Complex template causing stack overflow
+template<typename... Args>
+struct TypeList {
+template<template<typename...> class F>
+using apply = F<Args...>;
+// Chronos missed: recursive instantiation depth
+// Required: SFINAE and constexpr evaluation
+Listing 2: C++ template metaprogramming failure
+Root Cause: Template metaprogramming requires compile-time reasoning that exceeds current model capabilities.
+IX-B3 Case 3: Machine Learning Pipeline Data Leakage
+Bug Description: Validation accuracy drops 40% in production despite 95% test accuracy
+Failure Mode: Chronos identified standard preprocessing steps but missed subtle data leakage through feature normalization before train/test split.
+Root Cause: Requires understanding of statistical independence and temporal data dependencies in ML pipelines.
+IX-C Failure Pattern Analysis
+11122233344455566677788899910101005050501001001002.35.18.715.223.638.952.171.384.793.2Bug Complexity (1-10 scale)Failure Rate (%)
+Figure 20: Failure rate increases exponentially with bug complexity. The system struggles with highly complex bugs (complexity 8+) that require cross-module understanding.
+IX-D Mitigation Strategies and Future Directions
+1. Enhanced Retrieval for Edge Cases:
+  • Implement fallback to exhaustive search when confidence <0.3<0.3< 0.3
+  • Add cross-repository pattern matching for rare bugs
+  • Develop specialized retrievers for distributed/concurrent code
+2. Improved Understanding through Neuro-Symbolic Integration:
+  • Integrate theorem provers for correctness verification
+  • Add symbolic execution for path exploration
+  • Incorporate domain-specific reasoning modules
+3. Generation Safety Mechanisms:
+  • Mandatory test generation before fix application
+  • Rollback mechanisms for breaking changes
+  • Human-in-the-loop validation for critical systems
+The Bottom Line Despite these edge cases, Chronos achieves 87.1% debugging success compared to 22.9% for the next best system.
+Even in failure cases, Chronos provides actionable insights that accelerate manual debugging by 2.3x on average.
+X LIMITATIONS AND FUTURE WORK
+X-A Technical Limitations
+While Chronos advances autonomous debugging capabilities, several technical constraints remain:
+TABLE XXXV: Summary of Technical Limitations
+Limitation Impact Affected Scenarios Mitigation
+Extreme-scale latency 5s+ retrieval time Repos $>10M$ LOC Parallel expansion
+Memory cold start 23% lower success New projects $<1K$ commits Transfer learning
+Non-determinism 18.3% variance Distributed systems Deterministic replay
+Reasoning opacity 10GB traces/bug All debugging sessions Selective logging
+Memory coherence 2.1% conflicts Concurrent instances Eventual consistency
+Dynamic languages 41.2% accuracy Python/Ruby/JS Runtime instrumentation
+External dependencies 31% lower success API/DB bugs Mock generation
+Key Technical Constraints:
+  • Extreme-Scale Context Latency: O(k²n) complexity causes 5s+ retrieval times in 10M+ LOC repositories.
+  • Memory Cold Start: 23% lower success on projects with $<1K$ commits until memory builds over 2-3 weeks.
+  • Non-Determinism: 18.3% variance in distributed systems due to timing and resource contention.
+  • Reasoning Transparency: 10GB trace data per bug makes full interpretability infeasible.
+  • Dynamic Languages: 41.2% accuracy on Python/Ruby/JS due to runtime metaprogramming.
+  • External Dependencies: 31% lower success on API/database bugs without visibility into external states.
+X-A1 Failure Mode Analysis
+TABLE XXXVI: Common failure modes, frequencies, and mitigation strategies.
+Failure Mode Root Cause Frequency Mitigation Strategy
+Incomplete fixes Insufficient test coverage 12.3% Automated test generation
+Over-engineering Historical pattern bias 8.7% Confidence thresholding
+Context overflow Repository complexity 6.2% Hierarchical retrieval
+False positives Ambiguous error messages 4.8% Multi-source validation
+Regression introduction Side effect blindness 3.1% Impact analysis expansion
+X-B Research Directions: Extending the Debugging Paradigm
+Key research directions:
+  • Algorithmic Optimization: Sub-quadratic retrieval for 10M+ LOC repositories
+  • Visual Understanding: Screenshot analysis for UI/UX debugging
+  • Federated Learning: Cross-organization bug patterns with privacy
+  • Human-AI Collaboration: Interactive debugging with feedback loops
+  • Security Hardening: Defense against adversarial attacks
+X-C Deployment Architecture and Integration
 The proposed architecture extends beyond isolated debugging to comprehensive autonomous maintenance through a multi-tiered system design. The integration framework comprises:
@@
 Field trials with industry partners are ongoing to validate these findings at scale.
-VII-D Broader Impact
+X-D Broader Impact
 By enabling persistently self-healing and context-aware code maintenance, Chronos aims to shift an industry paradigm: reducing human toil and repetitive bug resolution, freeing engineers to focus on architecture, innovation, and user demands.
 As we scale deployment, it is crucial to steward responsible AI governance, data privacy, and an inclusive transition for developer workforces worldwide.
-VIII CONCLUSION
+DebugSuccessRoot CauseAccuracyRetrievalPrecisionMulti-FileUnderstandingFixQuality020202040404060606080808010010010071.289.283.578.9Evaluation MetricPerformance (%)ChronosClaude 4 OpusGPT-4.1Gemini 2.0 ProDeepSeek V3
+Figure 21: Comprehensive performance comparison across key debugging metrics. Chronos consistently outperforms state-of-the-art models by 3-5x across all evaluation dimensions.
+TABLE XXXVII: End-to-end debugging pipeline efficiency metrics
+System Avg. Time to Fix Iterations Token Usage Cost per Bug Success Rate
+Chronos 4.2 min 2.2 31.2K \$0.18 65.3%
+Claude 4 Opus + RAG 18.7 min 5.8 142.3K \$2.84 14.2%
+GPT-4.1 + LangChain 21.3 min 6.2 156.7K \$3.13 13.8%
+Gemini 2.0 Pro 19.5 min 5.5 134.8K \$1.35 13.4%
+Human Developer 35.8 min 3.4 N/A \$29.83 87.2%
+Based on average developer hourly rate of \$50
+XI CONCLUSION
 We have presented Chronos, a novel debugging-specific language model that addresses fundamental limitations in existing code understanding systems.
 Through specialized training on debugging workflows and a purpose-built architecture incorporating persistent memory and intelligent retrieval, Chronos demonstrates significant improvements over general-purpose language models in automated debugging tasks.
-Our comprehensive evaluation reveals that Chronos achieves a 65.3% success rate on real-world debugging benchmarks, representing a 6-7x improvement over state-of-the-art models including those with million-token contexts.
-This performance gain is attributed to three key technical contributions: (1) domain-specific pre-training on 15 million debugging instances including stack traces, fix commits, and CI/CD logs, (2) a persistent memory architecture that maintains cross-session knowledge of project-specific patterns, and (3) a hierarchical retrieval mechanism that enables effective reasoning over repository-scale contexts without computational constraints.
+Our comprehensive evaluation reveals that Chronos achieves a 65.3% success rate on real-world debugging benchmarks, representing a 4-5x improvement over the best 2025 models including Claude 4 Opus (14.2%), GPT-4.1 (13.8%), and Gemini 2.0 Pro (13.4%).
+This performance gain persists despite these models achieving state-of-the-art results on code generation tasks (Claude 4 Opus: 72.5% on SWE-bench).
+The disparity confirms our hypothesis: debugging is fundamentally different from code generation and requires specialized architectures.
+Key technical contributions enabling this breakthrough include: (1) domain-specific pre-training on 15 million debugging instances including stack traces, fix commits, and CI/CD logs, (2) Adaptive Graph-Guided Retrieval (AGR) that outperforms advanced RAG techniques like HyDE, Self-RAG, and FLARE by 2-3x on debugging tasks, (3) a persistent memory architecture that maintains cross-session knowledge, a capability absent in modern IDEs like Cursor and Windsurf, and (4) an autonomous debugging loop with iterative refinement based on test execution feedback.
 The implications of this work extend beyond immediate debugging applications.
 By demonstrating that specialized architectures and training regimes can dramatically improve performance on complex software engineering tasks, we provide evidence for the viability of task-specific language models in technical domains.
+While Chronos is built for debugging, its architecture naturally extends to long-term codebase governance , enabling intelligent memory of design decisions, bug patterns, and recovery heuristics.
 Future research directions include extending this approach to other software engineering workflows, investigating transfer learning between debugging domains, and exploring human-AI collaborative debugging frameworks.
@@
 ACKNOWLEDGMENTS
-Acknowledgments
 This work benefited from the feedback and real-world challenges shared by early-access engineering partners, enterprise pilot users, and the broader Kodezi community.